### v1

# Euclid Quick Data Release (Q1) point-like X-ray selected sources

Euclid Quick Data Release (Q1)
point-like X-ray selected sources
To better understand the role of active galactic nuclei (AGN) in galaxy evolution, it is crucial to achieve a complete and pure AGN census. X-ray surveys are key to this, but identifying their counterparts (CTPs) at other wavelengths remains challenging due to their larger positional uncertainties and limited availability of deeper, uniform ancillary data. Euclid is revolutionising this effort, offering extensive coverage of nearly the entire extragalactic sky, particularly in the near-infrared bands, where AGN are more easily detected. With the first Euclid Quick Data Release (Q1), we tested and validated methods for identifying, classifying, and determining the redshifts of Euclid CTPs to known point-like sources from major X-ray surveys, including XMM-Newton, Chandra, and eROSITA. Using Bayesian statistics, combined with machine learning (ML), as incorporated in the algorithm NWAY, we identify the CTPs to 11 286 X-ray sources from the three X-ray telescopes. For the large majority of 10 194 sources, the association is unique, with the remaining 10% of multi-CTP cases equally split between XMM-Newton and eROSITA. Six percent of the Euclid CTPs are detected in more than one X-ray survey. ML is then used to distinguish between Galactic (8%) and extragalactic (92%) sources. We computed photo-s using deep learning for the 8617 sources detected in the 10th data release of the DESI Legacy Survey, reaching an accuracy and a fraction of outliers of about 5%. Based on their X-ray luminosities, over 99% of CTPs identified as extragalactic are classified as AGN, most of which appear unobscured according to their hardness ratios. With this paper, we release our catalogue, which includes identifiers, basic X-ray properties, the reliability of the associations, and additional property extensions such as Galactic/extragalactic classifications and photometric/spectroscopic redshifts. We also provide probabilities for sub-selecting the sample based on purity and completeness, allowing users to tailor the sample according to their specific needs.
Key Words.:
Methods: statistical, data analysis; Surveys; Catalogues; Galaxies: active; X-rays: galaxies1 Introduction
Active galactic nuclei (AGN111Used interchangeably in both singular and plural cases), which rank among the most energetic phenomena in the Universe, play a pivotal role in shaping the evolutionary trajectories of galaxies throughout cosmic time. Their activity is fundamentally driven by the accretion of matter onto supermassive black holes (SMBHs) from their surrounding environment (e.g., Salpeter 1964; Lynden-Bell 1969; Pringle & Rees 1972; Shakura & Sunyaev 1973; Lynden-Bell & Pringle 1974; Kormendy & Ho 2013). Situated at the centres of galaxies, AGN emit intense radiation across the entire electromagnetic spectrum, ranging from radio waves to -rays (e.g., Elvis et al. 1994; Urry & Padovani 1995; Padovani et al. 2017). While this is particularly true for quasi-stellar objects (QSOs), it is important to note that the range of activity, emission, and interaction of AGN spans a broad spectrum, with different AGN exhibiting varying levels of energy outputs (e.g., Peterson 1997; Pierce et al. 2010; Pović et al. 2012; Bettoni et al. 2015).
Consequently, with enough coupling efficiency, this allows AGN to exert significant influence over both their host galaxies and the growth of their central SMBHs, where the interplay between AGN activity and the respective host is believed to regulate star formation and shape the morphology of galaxies. Here, energetic feedback drives powerful winds and outflows, impacting the surrounding interstellar and intergalactic medium (Greene et al. 2011; Cielo et al. 2018). These close connections are supported by fundamental scaling relations that link the SMBH mass to various galactic properties. Notable among these are the – relation, which ties BH mass to the velocity dispersion of stars in the galaxy’s bulge (Gebhardt et al. 2000; Ferrarese & Merritt 2000), and the – relation, which connects it to the stellar mass of the host galaxy (Magorrian et al. 1998; Häring & Rix 2004; Gültekin et al. 2009; Sani et al. 2011; Shankar et al. 2016; Suh et al. 2020), pointing to an intricate co-evolution of the systems (Kormendy & Ho 2013; Heckman & Best 2014; Madau & Dickinson 2014; Euclid Collaboration: Bisigello et al. 2024). However, the physical mechanisms driving this connection remain under debate, highlighting the need to construct a complete picture of AGN diversity and complexity to better understand galaxy formation and transformation across the history of the Universe (e.g., Aird et al. 2010; Buchner et al. 2015; Georgakakis et al. 2015; Ananna et al. 2017; Morganti 2017; Delvecchio et al. 2017; Husemann & Harrison 2018; Harrison & Ramos Almeida 2024).
Despite significant advancements over the years, the current AGN census remains incomplete due to inherent biases introduced to individual samples by various multi-wavelength selection methods, each favouring specific properties of the population (Messias et al. 2014; Padovani et al. 2017; Delvecchio et al. 2017; Lyu et al. 2022; Green et al. 2024). Comprehensive wavelength coverage is essential for identifying diverse AGN types across all redshifts, with each method capitalising on the dominant physical processes within its waveband to uncover unique facets of AGN activity, where the AGN dominates the spectral energy distribution (SED) of its host galaxy (Padovani et al. 2017). For instance, radio observations are effective in identifying jet-dominated AGN (De Breuck et al. 2002; Smolčić et al. 2017), while optical diagnostics (Feltre et al. 2016), such as emission line ratios (e.g., BPT diagrams as in Baldwin et al. 1981; Veilleux & Osterbrock 1987; Juneau et al. 2014), are widely used to distinguish AGN from star-forming galaxies by probing ionisation mechanisms. Other selections include colour or variability (e.g., Richards et al. 2002; Bongiorno et al. 2010; Bovy et al. 2012; Donley et al. 2012; Kirkpatrick et al. 2013; Peters et al. 2015; Palanque-Delabrouille et al. 2016; Lusso & Risaliti 2016). Soft X-ray selection excels at detecting unobscured accretion-powered emission (e.g., Hasinger 2008; Brandt & Hasinger 2005; Nandra et al. 2015) and can be complemented by mid-infrared (MIR) observations (e.g., Stern et al. 2012; Assef et al. 2013), which penetrate dense dust clouds effectively (Hickox & Alexander 2018). Notably, the significantly higher contrast between the intrinsic brightness of accreting SMBHs and their host galaxies in X-rays, as opposed to other wavelengths, makes this selection one of the most uncontaminated, providing pure AGN samples of larger diversity (Xue et al. 2011a; Donley et al. 2012). To combine different observational perspectives of AGN, these X-ray samples must then be paired with ancillary multi-wavelength data.
In the past, identifying multi-wavelength counterparts (CTPs) for X-ray selected sources was a significant hurdle in AGN characterisation and redshift estimation (Salvato et al. 2018). While combining data from multiple wavelengths is crucial for constructing a more complete picture of AGN, the process of cross-identifying sources remains complicated. Due to the significant positional uncertainties in X-ray surveys, confidently linking most X-ray sources to a single, definitive CTP in e.g., optical, infrared, or radio surveys is challenging, except for the brightest cases. These uncertainties, along with the uneven spatial coverage of ancillary data, further complicate the task of reliably pinpointing AGN across different surveys. Consequently, a simple coordinate match is insufficient, and more sophisticated methods are needed for accurate cross-identification. For this purpose, machine learning (ML) algorithms that integrate multi-wavelength source classification approaches have become increasingly prevalent (e.g., Cavuoti et al. 2014; Brescia et al. 2015; Karsten et al. 2023; Cooper et al. 2023; Daoutis et al. 2023; Zeraatgari et al. 2024; Pérez-Díaz et al. 2024; Mechbal et al. 2024).
While challenging to this date, the European Space Agency’s Euclid satellite will help overcome this limitation due to its uniform depth and wide coverage (Laureijs et al. 2011; Euclid Collaboration: Mellier et al. 2024). At the same time, Euclid’s depth results in very high source densities, increasing the risk of unforced misidentification and chance alignments. Over its six-year mission lifetime, Euclid will cover a total of 14 000 deg2 in the Euclid Wide Survey (EWS, Euclid Collaboration: Scaramella et al. 2022), expected to detect billions of sources. In addition to its cosmological objectives, Euclid will play a crucial role in the detection and characterization of AGN, expecting to detect at least 10 million of them (Euclid Collaboration: Bisigello et al. 2024; Euclid Collaboration: Lusso et al. 2024; Euclid Collaboration: Selwood et al. 2025). To achieve these goals, Euclid employs two main instruments: the VISible imager (VIS, Euclid Collaboration: Desprez et al. 2020; Euclid Collaboration: Cropper et al. 2024), covering wavelengths from 0.53 to 0.90 () and the Near-Infrared Spectrometer and Photometer (NISP, Maciaszek et al. 2022; Euclid Collaboration: Jahnke et al. 2024) operating in the near-infrared, with photometric measurements in three bands (, , and ), covering in total the wavelength range from 0.95 to 2.02 , while also including a slitless spectrograph for the detection of emission lines.
With the first Euclid Quick Data Release (Q1) we can now begin building and testing the necessary machinery for identifying and characterising AGN previously selected in the X-rays, where we aim to leverage Euclid’s depth, resolution and wavelength coverage to provide more secure associations. This initial testing will form the foundation for refining our methods and ensuring accurate AGN identification as more data becomes available over the next years.
In this paper, we address the task of reliably associating X-ray sources with their Euclid CTPs and investigate their respective NIR properties as a function of X-ray flux, among other properties. The paper is structured as follows: Sect. 2 gives an overview of the Q1 data. Section 3 introduces the X-ray source catalogues used in this work, while Sect. 4 outlines the approach to identifying CTPs. Section 5 describes how using a sample of sources clearly characterised with the 10th Data Release of the Legacy Survey (Dey et al. 2019), we assign a Galactic/extragalactic nature to our sources using Euclid photometry. In Sect. 6, we compute photometric redshifts for all CTPs. Section 7 displays CTP sample properties and Sect. 8 explains the release of the catalogues and how to use them. A summary and discussion are given in the final Sect. 9.
In this paper, unless stated otherwise, we express magnitudes in the AB system (Oke & Gunn 1983) and adopt a flat CDM cosmology with km s-1 Mpc-1, and , to facilitate direct comparison with similar works from literature.
2 Euclid Q1 fields
Euclid Q1 marks the initial data release with a footprint area of 63.1 deg2 (Euclid Collaboration: Mellier et al. 2024). It consists of observations making up the Euclid Deep Survey (EDS), selected for their considerable multi-wavelength coverage. However, the data is currently at the depth of the EWS, with a 5 depth of (AB)24.5 for point sources. The EDS targets three important Euclid Deep Fields (EDFs, see Fig. 1): the EDF Fornax (EDF-F), EDF South (EDF-S), and the EDF North (EDF-N). For more details, we refer the reader to Euclid Collaboration: Aussel et al. (2025).
The catalogues provide comprehensive photometric information across all four Euclid bands (Euclid Collaboration: McCracken et al. 2025; Euclid Collaboration: Polenta et al. 2025; Euclid Collaboration: Romelli et al. 2025). In addition to the Euclid bands, the catalogue is supplemented with ground-based photometry in the optical bands, where available, sourced from various instruments depending on their access to the northern or southern extragalactic sky. Photometric measurements are provided as both template/Sérsic model fits and fluxes extracted within apertures corresponding to radii of 1 – 4 times the seeing full-width at half maximum (FWHM).
3 The X-ray surveys
X-ray surveys are often constrained by the flux detection limit of the most sensitive X-ray telescopes, which typically operate with small fields of view. Therefore, surveys with different observational strategies need to be combined to facilitate the characterisation of the full AGN population. Pencil-beam surveys that provide exceptionally deep observations, valuable for probing faint AGN and for characterising the low-luminosity end of the AGN population, are limited in sky coverage and, consequently, miss rare (bright or distant) sources (e.g., Brusa et al. 2010; Hsu et al. 2014; Nandra et al. 2015; Marchesi et al. 2016; Luo et al. 2017; Oh et al. 2018). Subsequently, the data releases of XMM-Newton 4XMM DR14 (Webb et al. 2020) and Chandra Source Catalogue 2 (CSC 2.0, Evans et al. 2024) feature outstanding sensitivity for X-ray selected sources but cover relatively small regions of the sky.
On the other hand, wide-area surveys are crucial for complementing pointed observations and uncover those AGN missed (e.g., Ananna et al. 2017; Fotopoulou et al. 2016). Therefore, we incorporate data from the extended Roentgen Survey with an Imaging Array (eROSITA; Predehl et al. 2021) DR1 Main sample (Merloni et al. 2024). As part of the ongoing eROSITA All-Sky Survey (eRASS), these data provide homogeneous, soft X-ray coverage of the entire extragalactic sky. While eROSITA’s positional accuracy is somewhat pourer than that of XMM-Newton and Chandra, its unparalleled breadth allows us to include a significantly larger sample of sources, enhancing the statistical power of our analysis and ensuring a more comprehensive AGN census (see Fig. 1 and Footnote 2). In the following subsections, we describe the selection criteria applied to the catalogues from each X-ray survey.
X-ray catalogue | Number of sources | AND point-like | AND POSERR ¡ limit | EDF-F | EDF-N | EDF-S |
(1) | (2) | (3) | (4) | (5) | (6) | (7) |
4XMM DR14 | 692 109 | 612 521 | 612 504 | 4523 | 445 | 273 |
CSC 2.0 | 317 167 | 296 473 | 296 162 | 1016 | 506 | 147 |
eROSITA DR1 main | 898 812 | 879 153 | 878 984 | 1012 | 0 | 3369 |
3.1 XMM-Newton
The XMM-Newton space-telescope provides some of the deepest and most detailed X-ray imaging available, with a spatial resolution of approximately 6″ and positional uncertainties typically below 1 – 2″. The data used in this study include both targeted and serendipitous observations, selected from the 4XMM-DR14 catalogue333catalogue and data model description available at https://cdsarc.cds.unistra.fr/viz-bin/cat/IX/69, which offers multi-band flux measurements and quality flags. To prepare the sample, point-like sources were filtered based on their extent likelihood and positional uncertainties: SC_EXT_ML 10, SC_EXTENT 10 and SC_POSERR 10″, where the threshold for the positional uncertainty is motivated by the respective distribution as a function of flux as shown in Fig. 2. This effectively reduces the computational load for cross-matching with Euclid. To be consistent with other X-ray surveys, we compute the X-ray flux for the 0.5 – 2 keV band as
and its uncertainty as
by combining the 0.5 – 1 keV, and 1 – 2 keV band fluxes and errors.
3.2 Chandra
The Chandra X-ray space-telescope provides sub-arcsecond angular resolution, which allows precise localisation of sources and makes it particularly effective in crowded or complex fields. Our sample incorporates sources from the CSC 2.0 catalogue444https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=IX/57&-to=3. To subsample only point-like sources, we select sources with the extended flag, , set to zero. The 0.5 – 2 keV X-ray flux and its uncertainty are computed as
with Flux and Flux referring to the default CSC 2.0 soft and medium bands, respectively, while (B) and (b) represent the upper and lower 1 error margins. The positional uncertainties are determined using the 2 errors along the semi-major (0) and semi-minor (1) axes of the error ellipse provided in the source catalogue, as
and downsample the CSC 2.0 catalogue such that C_POSERR 10″, to be consistent with 4XMM DR14 (see Fig. 2).
3.3 eROSITA
The eROSITA-DE DR1 main catalogue555https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/Catalogues_dr1/ offers comprehensive coverage of the western galactic hemisphere. While it does not include the EDF-N region (see Fig. 1), with a half energy width (HEW) of 26″ and a soft X-ray energy range of 0.2 – 2.3 keV, eROSITA offers a uniform data set ideal for large statistical studies. The data preparation for eROSITA included filtering for reliable point-source detections with EXT_LIKE 0. Additionally, we required that all nine quality flags (QFs) are set to zero, and applied a cut in positional uncertainty of POS_ERR 20″(refer to Footnote 2).
The choice to permit larger positional uncertainties in the eROSITA DR1 sample, as opposed to the XMM-Newton and Chandra samples, is guided by the comparison of their error distributions shown in Fig. 2. This approach ensures the retention of the majority of the sample while excluding only a small number of outliers with exceptionally high positional errors. We intentionally avoid applying more restrictive criteria, such as minimum flux, signal-to-noise ratio (S/N), isolated environments, detection likelihoods, or other observational limitations, in order to prevent excluding sources that could potentially be identified by Euclid. An overview of the sample sizes is given in Footnote 2. By setting this threshold, we aim to optimise the balance between maintaining a robust sample and minimising the computational effort required for CTP identification within the defined error limits.
4 Identification of CTPs
Feature | Description |
dered_*_flux | deredened template flux in , , , and |
colour_* | deredened template colour for , , , , , and |
snr_* | S/N for template , , , and |
point_like_prob | probability between 0 and 1 that the source is point-like |
vis_det | binary flag, where |
Until recently, the identification of CTPs was primarily performed using maximum likelihood matching (e.g., Sutherland & Saunders 1992). This approach considers the separation between sources in the primary catalogue (e.g., X-ray) and the secondary multiwavelength catalogue, the positional uncertainties of the primary sources, and the magnitude distribution of sources in the ancillary catalogues within a certain radius from the primary sources. While this method has been applied successfully in cases with small positional errors and bright sources (see e.g., Naylor et al. 2013; Brusa et al. 2010; Xue et al. 2011b), its reliability decreases in situations where positional errors are large, and more than one photometric point is necessary for identifying the correct CTP or when multiple potential CTPs exist within the search radius. In such cases, positional matching often fails to provide a robust identification.
4.1 NWAY
To address these challenges, bayesian statistic needs to be invoked (Budavári & Szalay 2008). With this in mind, algorithms such as NWAY (Salvato et al. 2018), and XMATCH (Pineau et al. 2017) were developed. Unlike the others, NWAY also allows for the adoption of priors in addition to positional matches between multiple catalogues. It refines its matching procedure by being able to supplement one or more features, such as colours, magnitudes, and S/N, into likelihood ratios used for CTP identification. These priors, based on (anti-) correlations between source properties of the target and field populations, allow NWAY to improve the likelihood of finding true matches, boosting the accuracy significantly over methods that rely solely on positional information.
4.1.1 The integration of priors in NWAY
Colour and magnitude priors are posteriors that encode knowledge from data or probabilistic distributions that describe the expected properties of true CTPs for the sources in the primary catalogue. For example, an AGN selected in X-rays is likely to have distinct colours (e.g., redder MIR colours due to dust) and magnitudes compared to stars or inactive galaxies. The difference between the methods resides then in the adoption of specific features able to distinguish an X-ray emitter (regardless of its Galactic or extragalactic nature) from a random source in the field (Salvato et al. 2022). NWAY uses these priors to add additional dimensions to calculate a more nuanced likelihood ratio. This reduces ambiguity, especially in crowded fields or cases with large positional uncertainties.
For instance, in the ROSAT All-Sky Survey (2RXS, Boller et al. 2016), Salvato et al. (2018) improved CTP identification by incorporating a 2D prior based on the W2 magnitude and W1W2 colour from AllWISE (Wright et al. 2010), effectively separating out AGN. In the more recent application to the eROSITA Final Equatorial-Depth Survey (eFEDS, Brunner et al. 2022), Salvato et al. (2022) defined a prior that utilised not only magnitudes and colours from the Legacy Survey Data Release 10 (Dey et al. 2019), but also S/N in all bands (griz, W1, W2, W3, and W4), as well as properties, such as proper motion, from Gaia DR3 (Vallenari et al. 2023) when available. For further details on the formalism, we refer to the NWAY documentation777https://github.com/JohannesBuchner/NWAY.
4.1.2 Q1 prior
In this work, we have adopted the same principle as described in the previous subsection, but restricted ourselves to using only photometric features from Euclid. To build priors based on multiple features, NWAY can incorporate probabilities derived from ML models, such as a random forest (RF) classifier (the sklearn implementation, Pedregosa et al. 2011). We start by utilising a sample of secure CTPs to X-ray sources from 4XMM DR11 (Traulsen et al. 2020) and CSC 2.0 for training the RF as presented by Salvato et al. (2022). NWAY is then applied to identify CTPs for these X-ray sources in the EDF catalogues. A search radius of three times the positional error threshold is set around each X-ray source to include all potential EDF matches within this radius. This large search radius ensures that widely separated CTPs are not missed, even for sources with the largest positional uncertainties. The sky coverage of each X-ray sample per EDF is calculated to account for overlapping search windows. Specifically, while the coverage for the X-ray sample is given by its multi-order coverage map (MOC) or pointing, the EDF coverage is derived as
where corresponds to the number of X-ray sources and is the overlapping area of adjacent search windows. Following the methodology outlined in Salvato et al. (2018), these coverage areas are used to compute the number densities, which inform two key probabilities: p_any (the probability that an X-ray source has a CTP) and p_i (the probability that each EDF source within the search radius is the correct CTP). NWAY identifies the most likely CTP for each X-ray source by selecting the EDF source with the highest p_i value. This CTP is considered the primary match. To assess the reliability of this identification, the p_any value is used as a confidence threshold. All sources identified as primary matches (match_flag == 1) with p_any ¿ 0.85, form the target sample. In addition, for every X-ray source, all other EDF sources within the search radius that are not the primary match are collected into a field population, provided p_any ¡ 0.1. This field population serves as a representation of non-X-ray sources, providing a comparison set for training the RF. To label the data sets, X-ray CTPs are assigned a target class of ’1’, while field objects are labelled as ‘0‘. The Euclid-only training features utilised for the RF classifier are outlined in Footnote 6. At present, photometry and colours from different apertures are not included as features, as the aperture sizes vary between sources despite sharing the same designation.
Future data releases may standardise aperture sizes across all objects, also enhancing the RF classifier’s ability to leverage the relationship between light profiles and X-ray emission. To evaluate model performance, 15% of the training data set (6000 field and 1730 target sources) is randomly set aside as a test set. The RF model is configured with 200 decision trees, with splits allowed if at least 25 samples remain in a branch. Per decision tree construction, 10 photometric features are considered and bootstrap sampling is applied to the training set for building the ensemble of trees. Since the data set is highly imbalanced, with field objects vastly outnumbering CTPs of X-ray sources, a weighting scheme is implemented to automatically adjust the contribution of each class during training (Salvato et al. 2022). This ensures the model effectively learns to differentiate between the two classes despite the imbalance. The classifier outputs a probability score X-ray, the predicted likelihood of a candidate CTP being X-ray emitting. is used for class prediction, where NWAY retains candidates with if their astrometric configuration strongly supports a match. Importantly, the probability is derived exclusively from photometric properties, independent of positional uncertainties, enabling it to be seamlessly integrated into NWAY’s Bayesian framework, complementing astrometric priors with photometric information.
We quantify the classification quality by comparing the input labels and predicted classes as: (i) true positive (TP, where label 1 and prediction 1); (ii) true negative (TN, where label 0 and prediction 0); (iii) false positive (FP, where label 0 and prediction 1); and (iv) false negative (FN, where label 1 and prediction 0). The model’s performance is summarised in the confusion matrix shown in Fig. 3, assessed by acquiring these values for the test set (see e.g., Fotopoulou & Paltani 2018). We define the following measures of quality for the classifier:
accuracy (TPTN)/(TPTNFPFN) ;
precision TP/(TPFP) ;
recall TP/(TPFN) ;
fall-out FP/(TNFP) .
The trained model achieves strong performance, with high accuracy of 89% and recall of 91%, indicating that most real X-ray emitters are correctly identified. The contamination rate remains low, with a fractional leakage of 11%. The fall-out fraction can, in principle, be reduced by including more field objects, though at the price of reduced recall.
To better understand the model’s decision-making process, we performed a detailed feature importance analysis. This revealed that while most features contribute marginally to the classification, a few stand out as particularly informative. Specifically, features related to the de-reddened band, including the flux, the S/N, and respective colour combinations, headed by , as well as the point-like probability, emerge as the most significant predictors. For a more comprehensive overview of feature contributions, we refer the reader to Appendix A.
X-ray catalogue | X-ray sources | CTPs | AND | AND | AND | w/ spec- | |
in LS10 | S/Nr ¿= 3 | ||||||
4XMM DR14 | 5241 | 5885 | 5239 | 3774 | 3711 | 5486 | 1122 |
CSC 2.0 | 1669 | 1712 | 1666 | 684 | 669 | 1625 | 442 |
eROSITA DR1 | 4381 | 5048 | 4381 | 4159 | 4141 | 4375 | 477 |
Total: | 11 291 | 12 645 | 11 286 | 8617 | 8521 | 11 486 | 2044 |
4.2 Adding the X-ray emitter prior to NWAY
The a-priori probability of an Euclid source to be an X-ray emitter, , is now incorporated as a prior into the NWAY framework to enhance our CTP identification process. To visualise the positive impact of the inclusion of this prior, we plot the distribution of p_any both before and after the inclusion of the X-ray emitter prior (see Fig. 4). Incorporating the prior shifts the distribution of p_any to higher values, making the identification of true CTPs more reliable. Notably, the inclusion of the prior also changes the respective CTP in about 30% of the cases, indicating that for those sources that are indeed not the CTP after adding the prior, the p_any value has been decreased.
In total, we associate 12 645 CTP candidates. Among these, 11 286 (about ) sources have , indicating that they have the highest probability p_i to be considered reliable CTPs. The remaining candidates, with , represent secondary but plausible matches that are not ruled out entirely. We find 1092 X-ray sources with multiple CTPs, the vast majority of which is made up by XMM-Newton and eROSITA with only Chandra providing the required positional accuracy for Euclid to have few sources with multiple plausible CTPs. While most of the multi-CTP cases are made up of two candidates, a few cases with CTPs are introduced, predominantly by eROSITA, given its larger positional uncertainty. Notably, only five X-ray sources within the Q1 footprint (see Footnote 2) lack an association entirely, possibly due to boundary effects where the true CTP falls outside the region covered by the EDFs.
With this catalogue, several scenarios emerge in terms of the relationships between CTPs and X-ray sources.
1.
Two X-ray entries from differing surveys are assigned to a single CTP. This scenario suggests but does not confirm the possibility of a single physical source being detected by two different X-ray surveys.
2.
Two X-ray entries from the same survey are assigned to a single CTP. Here, we can confidently state that two distinct physical sources are selected, though both are associated with the same Euclid ID.
3.
Cases where a single Euclid ID is linked to multiple entries from e.g., two or even all three X-ray surveys.
4.
Two X-ray entries where the same X-ray source, detected by multiple surveys, is matched to distinct Euclid CTPs.
Among the surveys, only eROSITA uniformly covers two of the three fields. In contrast, other surveys consist of pointed observations with differing depths and catalogues across distinct bands, meaning sources are not always detected at the same flux levels. As a result, the overlap between catalogues is complex and interpreting source matches becomes challenging. At this stage, gaps in coverage further limit the significance of overlap checks. Keeping this in mind we find 674 pairs which either correspond to case one or two, as well as 16 triplets referring to case three as shown in Fig. 5. We deliberately choose not to quantify case four, as there is currently no definitive way to resolve it with certainty. A breakdown of the CTP sample is given in Footnote 8.
4.3 Purity and completeness
To calibrate a reliable p_any cut-off, we perform the NWAY procedure (using the same prior) on a duplicate sample of the X-ray sources, referred to as ’randoms‘. These randoms are generated by displacing the positions of the X-ray sources to lie beyond the search radius, ensuring that any matches with multiwavelength CTPs are purely coincidental. By comparing the distribution of p_any values for both the original X-ray sample and the randoms, we can assess the impact of the prior by examining the two histograms in Fig. 4. The figure shows how the p_any value for the randoms is usually lower than for the real sample of X-ray sources, indicating that in a random position in the sky, there are no sources that have the same features of a typical X-ray emitters. For the few sources for which p_any remains high, either the prior was not representative enough, or, deeper data in the future will potentially reveal a fainter X-ray source.
To further assess the impact of incorporating into the NWAY procedure, we compare the p_any values as a function of the separation distance between the primary (X-ray) and secondary (multi-wavelength) sources. Figure 6 presents the separation between the 4XMM sources and the associated CTP in EDF-F as a function of p_any. The sources with a higher probability of being correct also have a high probability of being X-ray emitters, even at larger separation, the latter being mostly smaller than 10″. CTPs of little separation and low p_any showcase scenarios of chance alignments where objects appear close but do not show significant . Next, we focus on evaluating the purity and completeness of the CTP samples based on how they are sub-selected. In this context, we define completeness as the cumulative fraction of sources with p_any higher than a certain value and purity the fraction of sources with p_any higher than the value in the random sample. This gives a measure of how often the CTP in the real sample is associated to a CTP only by chance.
Figure 7 shows both purity and completeness for each EDF and X-ray survey as well as the ’sweet spot’ for the best trade off between these quantities as a function of p_any. Variations in the distributions result from differences in the positional uncertainties of the X-ray surveys and the source densities within both the X-ray data and the corresponding EDF. Depending on the scientific goal, a user interested in creating a catalogue that is extremely complete or pure, is hence free to select sources with a p_any beyond a certain threshold. For the rest of the work, we adopt the intersection between the purity and completeness curves as the threshold. In all fields, purity and completeness are at least 80%, with the Chandra catalogues unsurprisingly having the most reliable associations, reaching values of about 90%. Interestingly, the impact of the higher spatial resolution of Chandra is clear in the EDF-F field, where Fornax, the second richest galaxy cluster of the local Universe is located. In such a dense environment, it is more challenging to identify the correct CTP given the resolution of 4XMM and eROSITA.
To conclude this part, we assess the stability in the method by examining the impact of RF hyper-parameters like minimum leaf size and features per tree. Variations in these parameters mainly shift the intersection of purity and completeness curves along the p_any axis, with smaller leaf sizes enhancing subclass specificity and subtly altering the intersection point.
4.4 Separation and magnitude distribution of CTPs
The distribution of observed X-ray-optical separations, normalised by the X-ray positional uncertainty, is presented in Fig. 8 as a function of the CTP’s -band magnitude. For a quality-selected sample of p_any 0.2, this distribution well aligns with the expected Rayleigh distribution for a scale factor of , indicating consistent behaviour with statistical expectations, demonstrating the reliability of the matching process (e.g., Salvato et al. 2022). The bulk of the sources have a mean separation of 1.″22 and = 20.65.
5 Characterisation and classification of CTPs
Once the CTPs have been identified, it is essential to categorise the sources in order to study the underlying physical processes and populations. The primary distinction is between extragalactic sources (such as galaxies or faint and bright AGN) and galactic sources (including stars and compact objects). In the following section, we outline the methodology used for classifying the sources and the validation tests conducted to ensure the accuracy of the classification, the results of which are summarised in Footnote 8.
5.1 Galactic and extragalactic sources
The four Euclid bands alone lack the discriminatory power required to reliably separate Galactic from extragalactic sources, given the broadness of the optical filter, where most of the differences within the SED occur. As demonstrated in Euclid Collaboration: Bisigello et al. (2024), no combination of Euclid-only photometry achieves satisfactory fractions of purity and completeness for this task. Moreover, the ground-based photometry added to the Q1 catalogues is not completely cross-calibrated at this stage, and differences in quality and bands available differ from field to field. For this reason, we decided to use the photometry available in LS10, providing us with well-calibrated and reasonable depth in from DECam Legacy Survey observations, including data from the Dark Energy Survey (DES, Dark Energy Survey Collaboration et al. 2016), Beijing-Arizona Sky Survey (BASS, Zou et al. 2017), and the Mayall -band Legacy Survey (MzLS, Silva et al. 2016), as well as W[1,2,3,4] from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE, Mainzer et al. 2011; Lang 2014; Meisner et al. 2017). We match the 11 286 CTPs to LS10 using a positional tolerance and find an association for 68% of our sources, see Footnote 8. Next, we apply a cut of S/N ¿ 3 in the , , , and W1 bands to minimise the relative impact of photometric uncertainties. The selected sources are then plotted on the W1 versus colour-colour plane, to distinguish between Galactic and extragalactic objects (see Fig. 9). Using the separation line defined in Salvato et al. (2022), we preliminarily classify sources based on their location relative to the threshold. Objects above the threshold are flagged as extragalactic, while those below are classified as Galactic.
These classifications are then used to create individual subplots (lower panels of Fig. 9), displaying the sources by their classification using only Euclid colours. As highlighted in Euclid Collaboration: Bisigello et al. (2024), the Galactic and extragalactic sources trace almost the same feature space, confirming that Euclid photometry alone, comprising the , , , and bands, does not provide sufficient information to reliably distinguish between the two populations.
5.2 Probabilistic classifier approach
To overcome this limitation, we follow the methodology outlined in Sect. 4.1.1, training an RF classifier using the same Euclid-only features listed in Footnote 6. However, this time we label the sample based on the classification derived from the versus plot, setting Galactic objects (396) to ‘1‘ and extragalactic objects (1000) to ‘0‘. Using the trained model, we assign a Galactic or extragalactic label to each Euclid CTP, providing a probabilistic classification for robust downstream analyses.
We present the respective confusion matrix in Fig. 10 and note a recall of 92%, with a limited fall-out of 4%. For more details on the feature correlations see Appendix A. To further validate the RF performance, we examine the galactic probability distribution () for the unique CTP candidates sample. Less than 10% of sources have , indicating effective separation of Galactic and extragalactic populations. A comparison with the Euclid Collaboration: Matamoro Zatarain et al. (2025) classification reveals that 97% of sources classified as Galactic there, also have , confirming good agreement. Additionally, the colour-coding of Fig. 11 reveals that most sources classified as being Galactic make up the bright end of the sample, with the most intense colour gradient surrounding and a couple of sources towards high interestingly appearing increasingly fainter again, probably due to different populations of stars, subject to their location in the extragalactic sky.
6 Redshifts
The catalogues of CTPs we make available in this paper include spectroscopic redshifts (spec-s), where available, by matching their coordinates to a compilation of publicly available redshifts (Kluge et al. 2024). However, this compilation is rich in duplications, hence the details of the cleaning are described in Sect. 3.1 of Saxena et al. (2024). While spec-s provide highly accurate distance measurements, they are inherently limited to a subset of typically brighter objects. As a result only 16% of the CTP sample benefits from publicly available spec- coverage (refer to Footnote 8).
6.1 Photometric redshifts
To achieve a more complete characterization of the CTPs, photometric redshifts (photo-s) are essential. However, achieving reliable photo-s for AGN remains significantly more challenging than for inactive galaxies, despite notable advances in recent years (see Salvato et al. 2019, for a review).
The primary difficulty lies in the overwhelming emission produced by the accretion onto the SMBH, which often masks the spectral features of the host galaxy, crucial for breaking colour-redshift degeneracies. The current version of photo-s released for Q1 assumes the sources to be inactive galaxies and does not account for the possible contribution from an active nucleus. The absence of AGN SEDs in the template library makes it difficult to accurately constrain their contributions using photometry alone, resulting in miscalculations of the physical parameters for each source derived indirectly. Instead, studies have demonstrated that incorporating aperture or pixel-based analyses can significantly enhance the accuracy of such photo- estimates. Consequently, to compute reliable photo-s for AGN-dominated objects in a dedicated effort, we utilise PICZL (Roster et al. 2024), an ML algorithm that significantly improves photo- estimation for AGN by directly predicting redshifts from imaging data, eliminating the need for manual feature extraction or additional surveys. Building on its predecessor, Circlez (Saxena et al. 2024), which analysed aperture-based photometric variations, PICZL advances to pixel-level resolution, leveraging the raw images for improved precision. As a result, for sources bright enough to successfully be matched to the LS10 catalogue, we are able to produce reliable photo-s (refer to Footnote 8). This is possible since the training sample constructed for PICZL is formed mostly by X-ray selected AGN detected in LS10, enabling us to use the same model. As in the original work, the uncertainty of the photo-s in terms of 1 error, increases with the faintness of the sources (see Roster et al. 2024, for more details).
For all EDF CTPs of an X-ray source detected in LS10, PICZL provides full redshift posterior distributions (PDZ) derived from de-reddened calibrated images in the , , , and bands, supplemented by , , , and catalogue-based aperture photometry without imposing any S/N cuts. The method is agnostic to the morphological classification of the source, applying equally to extended and point-like objects. As output, PICZL provides the dominant mode of the PDZ as a point prediction, along with uncertainties quantified as 1 and 3 confidence intervals. Notably, we compute photo-s also for those CTPs matched to LS10, which we believe to be Galactic. This approach accounts for the inherent uncertainty in the probabilities produced by our model and classification threshold. To ensure flexibility for users, we include both classifications and photo- estimates in the released catalogue (see Sect. 8). This allows users to tailor their selection criteria to their preferences while retaining access to the full set of photo- values.
6.2 Photo-z quality
To evaluate the reliability of our photo- estimates, we employ standard metrics commonly implemented in the literature, including (a) the accuracy as a measure of the scatter between the prediction and truth values of the sample, defined as
and (b) the fraction of outliers, , which quantifies the proportion of sources whose redshift estimates deviates significantly from their corresponding spectroscopic redshift. Specifically, this is defined as the fraction of objects for which
We evaluated the performance of the photo-s by comparing them to the spec-s, as shown in Fig. 12. For the subsample of sources with ¡ 21.5, which corresponds to the brighter and more reliable sources in our data set, we find that the fraction of outliers is approximately 5%, with a of roughly 0.04 as opposed to and observed for the entire sample. Additionally, the catastrophic outliers at are faint, galaxy-dominated objects ( ¿ 24) with large photometric errors in LS10. In scientific studies involving photometric redshifts, it is crucial to account for such uncertainties and in turn, photometric redshift algorithms should therefore make sure to provide reliable errors. PICZL achieves this effectively as discussed in detail in Roster et al. (2024). In the right panel of Fig. 12 this is evidenced by the smaller errors closely tracing the identity line, while larger errors appropriately account for sources of uncertainty.
In Fig. 13, we present the cumulative distribution of CTP band magnitudes along with the respective fractions of sources with spectroscopic and/or and photometric redshifts. The figure reveals that the sample is almost complete in redshift up to around . For fainter sources, photo- estimation is either not possible due to the limited number of photometric data points and significant errors, or the source, though clearly detected in Euclid, is not visible in LS10 due to its shallower depth. In the future, when we have upgraded PICZL to be applied to Euclid images and have integrated data from the Legacy Survey of Space and Time (LSST; Ivezic et al. 2019), photometric redshifts for AGN in Euclid will dramatically improve.
7 Properties of the CTP sample
Having successfully selected, classified, and vetted our CTPs, we now turn to an analysis of their properties. From this point forward, we refer to CTPs as those sources meeting the criteria of and p_any. This refined selection yields a total of 9294 sources from the initial 11 286 entries having match_flag=1.
Figure 14 illustrates the magnitude distribution of our CTPs, grouped by the match between EDFs and the corresponding X-ray survey (4XMM, CSC2, and eROSITA DR1, respectively). The 4XMM and CSC 2.0 distributions show their ability to detect fainter X-ray sources, which correlates with the identification of fainter optical CTPs. In contrast, the eROSITA DR1 distributions exhibit brighter median magnitudes, consistent with the shallowness of the survey (see Fig. 2). Additionally, the figure suggests a possible bimodal distribution at the faint end in the optical band for both 4XMM and CSC 2.0, at least in EDF-N and EDF-S. The majority of ources in the second, faint peak of the distribution show X-ray fluxes , with no obvious drop off in p_any, and the distribution will be studied further in other papers.
7.1 Colour-redshift relation
Redshifts, spectroscopic and photometric, are available for 79% of the CTPs, allowing us to analyse their Euclid colours as function of redshift, as shown in Fig. 15. The colour-coded kernel density is plotted along the y-axis dimension within each bin, providing a visual trace of how the physical colours of the CTP sample vary as a function of redshift.
A notable feature in the data is the detection of the Balmer break at , particularly prominent in colours including the filter. Unsurprisingly, NIR-only colour combinations display relatively few discernible features (compare e.g., figure 7 in Temple et al. 2021). As iterated before, the Euclid filters alone lack the fidelity to catch variations in the SED of sources, a caveat that we will be able to overcome in combination with upcoming surveys, especially LSST.
7.2 X-ray luminosities
For the rest of the analysis, we restrict the CTP sample to those 7345 extragalactic sources with a soft X-ray flux S/N 2 in order to reduce potential spurious X-ray detection. In Fig. 16, we observe a clear distinction between pointed observations taken with XMM-Newton and Chandra at various depths versus the homogenous all-sky coverage provided by eROSITA. The complete coverage of eROSITA results in a tighter correlation between luminosity and redshift, despite the presence of a few points away from the general trend, which have low S/N (thus indicating potential spurious sources). Another reason for their offset from the general trend could be an incorrect photometric redshift. On the other hand, XMM-Newton and Chandra catalogues include a mixture of pointed observations resulting in a less smooth distribution between redshift and luminosity.
In all three panels, we indicate with open circle sources with low S/Nr, potentially corresponding to those highlighted red in the left panel of Fig. 12, representing examples of increased redshift uncertainties. Notably, most of the results shown, focus on objects with redshifts of . Sources with higher redshifts, while present, should be approached with caution due to potential inaccuracies. However, the future availability of LSST photometry will enable the exploration of this parameter regime by probing fainter sources with smaller errors, thus improving the reliability of redshift estimates for these populations.
In studies such as that of Georgakakis & Nandra (2011), AGN are typically defined as sources with X-ray luminosities exceeding , since normal galaxies may be an important source of contamination below this limit (Lehmer et al. 2012; Aird et al. 2015). However, it is important to note that this threshold was established based on pencil-beam surveys, suggesting that it should be revisited now that wide-area surveys and deep optical data are accessible. X-ray luminosities of AGN in the soft X-ray band (0.5 – 2 keV) can vary significantly, depending on the AGN type and redshift. In literature, values can range from or even less when considering low-luminosity AGN (Hasinger et al. 2005; Ebrero et al. 2009; Civano et al. 2012; Kawamuro et al. 2013). Applying the criterion of , we classify 6243 out of the 6295 sources with computed X-ray luminosities as AGN.
A more inclusive approach is presented in Georgakakis et al. (2007) for example, where X-ray detected starburst galaxy candidates, though more likely in case of Chandra as opposed to eROSITA (Kyritsis et al. 2025), can represent about 20% of the X-ray source population in the luminosity interval . This fraction is also likely to be an upper limit because some of the galaxy candidates may turn out to be low–luminosity AGN, such that distinguishing between these sources can be challenging. Using our sources from XMM-Newton, where both soft and hard (2 – 12 keV)999We assume that at zero order, the luminosities between (2 – 10) and (2 – 12) keV are approximately the same. band fluxes are available, we show that, to retain all sources in this parent catalogue that are classified as AGN using the hard band, a soft-band luminosity threshold of is required (refer to Fig. 17). Given the results of both all-sky and pencil beam surveys, at soft luminosities in the range , AGN still appear to represent the dominant population. Nonetheless, future studies are required to determine the nature of these low-luminosity AGN at soft energies. Applying this more inclusive threshold defined in Fig. 17 to our extragalactic CTP sample, 6292 of the 6295 sources that we were able to compute X-ray luminosities for, meet the AGN definition criterion.
7.3 AGN type
In Fig. 18, we plot the respective -band magnitudes as a function of the soft X-ray fluxes for all CTPs classified as extragalactic. We add the —X/O— = 1 lines (Maccacaro et al. 1988), which highlight the typical AGN locus (see also e.g., Salvato et al. 2011; Civano et al. 2012), defined as
where is the X-ray flux in a given energy range, and correspond to the magnitude and a constant which both depend on the specific filter used in the optical observations. For we find . To further assess the type of our sources, we colour-code the plots in the bottom row of Fig. 18 by the X-ray luminosity and the top row by hardness ratio (Park et al. 2006; Cappelluti et al. 2007; Salvato et al. 2008), defined as
where a negative HR value indicates (type I) unobscured sources, whereas a positive HR value suggests harder, likely obscured (type II) sources. Therefore, we include fluxes from the hard X-ray band where available. Specifically, we consider fluxes from the 2.0 – 4.5 keV range for 4XMM DR14, the 2.0 – 7.0 keV range for Chandra, and the 2.3 – 5.0 keV energy range for eROSITA DR1. This ensures as much consistency as possible across the different surveys, allowing us to utilise similar energy bands for comparison and analysis.
Since the markers in Fig. 18 are plotted in order of increasing hardness ratio, many sources in the background are being covered up and predominantly fall within the X/O boundaries, classifying them as bright, unobscured type I sources, with different populations visible for every X-ray survey. This trend is most pronounced for eROSITA, where the distribution is more concentrated within the boundaries. A similar pattern is observed for XMM-Newton, though it includes a population of very soft sources lying below the X/O boundaries, potentially corresponding to starburst galaxies. For Chandra, the general trend also holds; however, there is an additional presence of very hard sources overlapping with very soft ones below the X/O boundary, particularly in the bright optical and faint X-ray regime. The majority of these outliers are sources with S/N 3 in the soft X-ray band, suggesting that at least some of them may be spurious detections. It is also plausible that these sources are intrinsically variable, suggesting that they were observed in different states of appearance. Overall, our distribution follows the findings of Merloni et al. (2014), which show that only approximately 80% of AGN selected by their SED, spectral features and HR, match their expected type I/II class.
8 Release of the catalogue
The catalogue containing the CTPs to the point-like X-ray sources in the EDFs and their properties are available through CDS/Vizier101010https://vizier.cds.unistra.fr. A detailed description of the columns and their content is provided in Appendix B. The catalogue includes unique identifiers for all relevant surveys as well as the basic X-ray and Euclid properties (columns 1 – 13). For the full list of available columns, please refer to the original catalogues introduced in Sect. 3 and e.g., Euclid Collaboration: McCracken et al. (2025); Euclid Collaboration: Polenta et al. (2025); Euclid Collaboration: Romelli et al. (2025). Columns 14 – 24 present the results of the CTP association, followed by key parameters derived from NWAY. Photometric data from LS10 are reported in columns 25 – 28. Spectroscopic information, when available, is included in columns 29 – 33. Lastly, the photometric redshift parameters from the PICZL algorithm, as well as quantities computed with such, are listed in columns 34 – 40.
9 Conclusions and outlook
In this paper we have presented the procedure adopted for the identification, classification, and study of the Euclid CTPs to the point-like sources selected from XMM-Newton, Chandra, and eROSITA X-ray, detected in Q1. We summarise the most important results in the following.
Utilising NWAY (Salvato et al. 2018), supplemented by a prior on X-ray emission, based on properties of a 4XMM DR11, and CSC 2.0 training sample, we achieved highly accurate CTP identifications relying on Euclid photometry only (see Sect. 4). Randomising the X-ray position and repeating the same procedure for the association, we quantify for each X-ray source the probability of chance associations. Users will be able to create subsamples of CTPs of desired purity and completeness by applying more restrictive or inclusive p_any (the probability of an X-ray source having a CTP) cuts (see Fig. 7).
Altogether, we identify 12 645 CTPs, with only about 10% of the sources having more than a single possible CTP, mostly in the 4XMM and eROSITA surveys, due to their larger positional uncertainties, compared to Chandra. We attribute this performance to the availability of extensive X-ray detected samples with secure CTPs, and deep, homogenised multiwavelength photometry spanning optical to MIR wavelengths, enabling robust SED construction for training. Further enhancements in expanding to broader multiwavelength coverage will improve the quality of CTP identification, particularly with Euclid DR1 in mind.
We trained a RF algorithm to classify objects as Galactic or extragalactic using solely Euclid photometry in Sect. 5. We previously trained on a sample of secure Galactic/extragalactic sources detected in LS10, using a over W1 thresholds, as presented in Salvato et al. (2022), effectively breaking the degeneracy in classification that arises when relying on Euclid photometry alone (see Euclid Collaboration: Bisigello et al. 2024). We validate these classifications by comparing to Euclid Collaboration: Matamoro Zatarain et al. (2025).
We also added spectroscopic redshifts for 16% of the CTP sample using publicly available data. For the 76% of the sample detected in LS10, we computed photometric redshifts using PICZL (Roster et al. 2024), reaching up to , although at these high redshifts the values are associated with large uncertainties. With all these redshifts, we investigated the colour-redshift relation of our sample in addition to computing their X-ray luminosities (refer to Sect. 6). We find that almost all of our CTPs pass the threshold to be considered an AGN, with most of them appearing as soft, bright, and unobscured type I sources.
As the primary outcome of this paper, we present the CTP catalogue, which includes a wealth of information such as IDs, basic X-ray properties, NWAY outputs, classification probabilities, and redshift data (see Sect. 8). By providing probabilities for all applicable sample characteristics, we ensure users have the flexibility to make informed selections tailored to their specific scientific goals, without being constrained by predefined thresholds or assumptions.
The RF models used for the identification of the CTPs and their classification into Galatic/extragalactic, can be applied blindly to the entire Q1 catalogue. For each source, they provide the probability of being an X-ray emitter, as well as an extragalactic source, despite the fact of not having been detected by any of the three X-ray surveys yet. Consequently, as illustrated in Figure Fig. 19, limiting the selection of Q1 sources to those with and , while excluding Euclid sources classified as CTPs in this study, yields a final sample of 135 303 objects. Notably, approximately 400 of these have been successfully cross-matched geometrically with the little red dots sample introduced by Euclid Collaboration: Bisigello et al. (2025), predominantly occupying the bright end of the magnitude distribution. A forthcoming paper will explore the multi-wavelength nature of likely extragalactic high sources that remain undetected in current X-ray surveys.
Acknowledgements.
[…]
- Aird et al. (2015) Aird, J., Coil, A. L., Georgakakis, A., et al. 2015, MNRAS, 451, 1892
- Aird et al. (2010) Aird, J., Nandra, K., Laird, E. S., et al. 2010, MNRAS, 401, 2531
- Ananna et al. (2017) Ananna, T. T., Salvato, M., LaMassa, S., et al. 2017, ApJ, 850, 66
- Assef et al. (2013) Assef, R. J., Stern, D., Kochanek, C. S., et al. 2013, ApJ, 772, 26
- Baldwin et al. (1981) Baldwin, J. A., Phillips, M. M., & Terlevich, R. 1981, PASP, 93, 5
- Bettoni et al. (2015) Bettoni, D., Falomo, R., Kotilainen, J. K., Karhunen, K., & Uslenghi, M. 2015, MNRAS, 454, 4103
- Boller et al. (2016) Boller, T., Freyberg, M. J., Trümper, J., et al. 2016, A&A, 588, A103
- Bongiorno et al. (2010) Bongiorno, A., Mignoli, M., Zamorani, G., et al. 2010, A&A, 510, A56
- Bovy et al. (2012) Bovy, J., Myers, A. D., Hennawi, J. F., et al. 2012, ApJ, 749, 41
- Brandt & Hasinger (2005) Brandt, W. N. & Hasinger, G. 2005, ARA&A, 43, 827
- Brescia et al. (2015) Brescia, M., Cavuoti, S., & Longo, G. 2015, MNRAS, 450, 3893
- Brunner et al. (2022) Brunner, H., Liu, T., Lamer, G., et al. 2022, A&A, 661, A1
- Brusa et al. (2010) Brusa, M., Civano, F., Comastri, A., et al. 2010, ApJ, 716, 348
- Buchner et al. (2015) Buchner, J., Georgakakis, A., Nandra, K., et al. 2015, ApJ, 802, 89
- Budavári & Szalay (2008) Budavári, T. & Szalay, A. S. 2008, ApJ, 679, 301
- Cappelluti et al. (2007) Cappelluti, N., Hasinger, G., Brusa, M., et al. 2007, ApJS, 172, 341
- Cavuoti et al. (2014) Cavuoti, S., Brescia, M., D’Abrusco, R., Longo, G., & Paolillo, M. 2014, MNRAS, 437, 968
- Chen et al. (2022) Chen, H., Lundberg, S. M., & Lee, S.-I. 2022, Nature Communications, 13
- Cielo et al. (2018) Cielo, S., Bieri, R., Volonteri, M., Wagner, A. Y., & Dubois, Y. 2018, MNRAS, 477, 1336
- Civano et al. (2012) Civano, F., Elvis, M., Brusa, M., et al. 2012, ApJS, 201, 30
- Cooper et al. (2023) Cooper, N., Dainotti, M. G., Narendra, A., Liodakis, I., & Bogdan, M. 2023, MNRAS, 525, 1731
- Cuillandre et al. (2024) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2024, A&A, accepted, arXiv:2405.13501
- Daoutis et al. (2023) Daoutis, C., Kyritsis, E., Kouroumpatzakis, K., & Zezas, A. 2023, A&A, 679, A76
- Dark Energy Survey Collaboration et al. (2016) Dark Energy Survey Collaboration, Abbott, T., Abdalla, F. B., et al. 2016, MNRAS, 460, 1270
- De Breuck et al. (2002) De Breuck, C., van Breugel, W., Röttgering, H., & Carilli, C. 2002, in ASP, Vol. 284, IAU Colloq. 184: AGN Surveys, ed. R. F. Green, E. Y. Khachikian, & D. B. Sanders, 275
- Delvecchio et al. (2017) Delvecchio, I., Smolčić, V., Zamorani, G., et al. 2017, A&A, 602, A3
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168
- Dey et al. (2022) Dey, B., Andrews, B. H., Newman, J. A., et al. 2022, MNRAS, 515, 5285
- Donley et al. (2012) Donley, J. L., Koekemoer, A. M., Brusa, M., et al. 2012, ApJ, 748, 142
- Ebrero et al. (2009) Ebrero, J., Carrera, F. J., Page, M. J., et al. 2009, A&A, 493, 55
- Elvis et al. (1994) Elvis, M., Wilkes, B. J., McDowell, J. C., et al. 1994, ApJS, 95, 1
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted
- Euclid Collaboration: Bisigello et al. (2024) Euclid Collaboration: Bisigello, L., Massimo, M., Tortora, C., et al. 2024, A&A, 691, A1
- Euclid Collaboration: Bisigello et al. (2025) Euclid Collaboration: Bisigello, L., Rodighiero, G., Fotopoulou, S., et al. 2025, A&A, submitted
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
- Euclid Collaboration: Desprez et al. (2020) Euclid Collaboration: Desprez, G., Paltani, S., Coupon, J., et al. 2020, A&A, 644, A31
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
- Euclid Collaboration: Lusso et al. (2024) Euclid Collaboration: Lusso, E., Fotopoulou, S., Selwood, M., et al. 2024, A&A, 685, A108
- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H., Benson, K., et al. 2025, A&A, submitted
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
- Euclid Collaboration: Polenta et al. (2025) Euclid Collaboration: Polenta, G., Frailis, M., Alavi, A., et al. 2025, A&A, submitted
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
- Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
- Euclid Collaboration: Selwood et al. (2025) Euclid Collaboration: Selwood, M., Fotopoulou, S., Bremer, M. N., et al. 2025, A&A, 693, A250
- Evans et al. (2024) Evans, I. N., Evans, J. D., Martínez-Galarza, J. R., et al. 2024, ApJS, 274, 22
- Feltre et al. (2016) Feltre, A., Charlot, S., & Gutkin, J. 2016, MNRAS, 456, 3354
- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
- Fotopoulou et al. (2016) Fotopoulou, S., Pacaud, F., Paltani, S., et al. 2016, A&A, 592, A5
- Fotopoulou & Paltani (2018) Fotopoulou, S. & Paltani, S. 2018, A&A, 619, A14
- Gebhardt et al. (2000) Gebhardt, K., Bender, R., Bower, G., et al. 2000, ApJ, 539, L13
- Georgakakis et al. (2015) Georgakakis, A., Aird, J., Buchner, J., et al. 2015, MNRAS, 453, 1946
- Georgakakis & Nandra (2011) Georgakakis, A. & Nandra, K. 2011, MNRAS, 414, 992
- Georgakakis et al. (2007) Georgakakis, A., Rowan-Robinson, M., Babbedge, T. S. R., & Georgantopoulos, I. 2007, MNRAS, 377, 203
- Green et al. (2024) Green, K., Elmer, E., Maltby, D. T., et al. 2024, MNRAS, 531, 2551
- Greene et al. (2011) Greene, J. E., Zakamska, N. L., Ho, L. C., & Barth, A. J. 2011, ApJ, 732, 9
- Gültekin et al. (2009) Gültekin, K., Richstone, D. O., Gebhardt, K., et al. 2009, ApJ, 698, 198
- Häring & Rix (2004) Häring, N. & Rix, H.-W. 2004, ApJ, 604, L89
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
- Hasinger (2008) Hasinger, G. 2008, A&A, 490, 905
- Hasinger et al. (2005) Hasinger, G., Miyaji, T., & Schmidt, M. 2005, A&A, 441, 417
- Heckman & Best (2014) Heckman, T. M. & Best, P. N. 2014, A&A, 52, 589
- Hickox & Alexander (2018) Hickox, R. C. & Alexander, D. M. 2018, ARA&A, 56, 625
- Hsu et al. (2014) Hsu, L.-T., Salvato, M., Nandra, K., et al. 2014, ApJ, 796, 60
- Husemann & Harrison (2018) Husemann, B. & Harrison, C. M. 2018, Nature Astronomy, 2, 196–197
- Ivezic et al. (2019) Ivezic, Z., Kahn, S. M., Tyson, J. A., et al. 2019, ApJ, 873, 111
- Juneau et al. (2014) Juneau, S., Bournaud, F., Charlot, S., et al. 2014, ApJ, 788, 88
- Karsten et al. (2023) Karsten, J., Wang, L., Margalef-Bentabol, B., et al. 2023, A&A, 675, A159
- Kawamuro et al. (2013) Kawamuro, T., Ueda, Y., Tazaki, F., & Terashima, Y. 2013, ApJ, 770, 157
- Kirkpatrick et al. (2013) Kirkpatrick, A., Pope, A., Charmandaris, V., et al. 2013, ApJ, 763, 123
- Kluge et al. (2024) Kluge, M., Comparat, J., Liu, A., et al. 2024, A&A, 688, A210
- Kormendy & Ho (2013) Kormendy, J. & Ho, L. C. 2013, ARA&A, 51, 511
- Kyritsis et al. (2025) Kyritsis, E., Zezas, A., Haberl, F., et al. 2025, A&A, 694, A128
- Lang (2014) Lang, D. 2014, ApJ, 147, 108
- Laureijs et al. (2011) Laureijs, R., Amiaux, J., Arduini, S., et al. 2011, ESA/SRE(2011)12, arXiv:1110.3193
- Lehmer et al. (2012) Lehmer, B. D., Xue, Y. Q., Brandt, W. N., et al. 2012, ApJ, 752, 46
- Lundberg & Lee (2017) Lundberg, S. & Lee, S.-I. 2017, arXiv e-prints, arXiv:1705.07874
- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
- Lusso & Risaliti (2016) Lusso, E. & Risaliti, G. 2016, ApJ, 819, 154
- Lynden-Bell (1969) Lynden-Bell, D. 1969, Nature, 223, 690
- Lynden-Bell & Pringle (1974) Lynden-Bell, D. & Pringle, J. E. 1974, MNRAS, 168, 603
- Lyu et al. (2022) Lyu, J., Alberts, S., Rieke, G. H., & Rujopakarn, W. 2022, ApJ, 941, 191
- Maccacaro et al. (1988) Maccacaro, T., Gioia, I. M., Wolter, A., Zamorani, G., & Stocke, J. T. 1988, ApJ, 326, 680
- Maciaszek et al. (2022) Maciaszek, T., Ealet, A., Gillard, W., et al. 2022, in SPIE, Vol. 12180, Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, ed. L. E. Coyle, S. Matsuura, & M. D. Perrin, arXiv:2210.10112
- Madau & Dickinson (2014) Madau, P. & Dickinson, M. 2014, ARA&A, 52, 415
- Magorrian et al. (1998) Magorrian, J., Tremaine, S., Richstone, D., et al. 1998, AJ, 115, 2285
- Mainzer et al. (2011) Mainzer, A., Bauer, J., Grav, T., et al. 2011, ApJ, 731, 53
- Marchesi et al. (2016) Marchesi, S., Lanzuisi, G., Civano, F., et al. 2016, ApJ, 830, 100
- Mechbal et al. (2024) Mechbal, S., Ackermann, M., & Kowalski, M. 2024, A&A, 685, A107
- Meisner et al. (2017) Meisner, A. M., Lang, D., & Schlegel, D. J. 2017, ApJ, 153, 38
- Merloni et al. (2014) Merloni, A., Bongiorno, A., Brusa, M., et al. 2014, MNRAS, 437, 3550
- Merloni et al. (2024) Merloni, A., Lamer, G., Liu, T., et al. 2024, A&A, 682, A34
- Messias et al. (2014) Messias, H., Afonso, J. M., Salvato, M., Mobasher, B., & Hopkins, A. M. 2014, A&A, 562, A144
- Morganti (2017) Morganti, R. 2017, Frontiers in Astronomy and Space Sciences, 4, 42
- Nandra et al. (2015) Nandra, K., Laird, E. S., Aird, J. A., et al. 2015, ApJS, 220, 10
- Naylor et al. (2013) Naylor, T., Broos, P. S., & Feigelson, E. D. 2013, ApJS, 209, 30
- Oh et al. (2018) Oh, K., Koss, M., Markwardt, C. B., et al. 2018, ApJS, 235, 4
- Oke & Gunn (1983) Oke, J. B. & Gunn, J. E. 1983, ApJ, 266, 713
- Padovani et al. (2017) Padovani, P., Alexander, D. M., Assef, R. J., et al. 2017, A&A Rev., 25, 2
- Palanque-Delabrouille et al. (2016) Palanque-Delabrouille, N., Magneville, C., Yèche, C., et al. 2016, A&A, 587, A41
- Park et al. (2006) Park, T., Kashyap, V. L., Siemiginowska, A., et al. 2006, ApJ, 652, 610
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Peters et al. (2015) Peters, C. M., Richards, G. T., Myers, A. D., et al. 2015, ApJ, 811, 95
- Peterson (1997) Peterson, B. M. 1997, An Introduction to Active Galactic Nuclei (Cambridge University Press)
- Pierce et al. (2010) Pierce, C. M., Lotz, J. M., Salim, S., et al. 2010, MNRAS, 408, 139
- Pineau et al. (2017) Pineau, F. X., Derriere, S., Motch, C., et al. 2017, A&A, 597, A89
- Pović et al. (2012) Pović , M., Sánchez-Portal, M., García, A. M. P., et al. 2012, A&A, 541, A118
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Pringle & Rees (1972) Pringle, J. E. & Rees, M. J. 1972, A&A, 21, 1
- Pérez-Díaz et al. (2024) Pérez-Díaz, V. S., Martínez-Galarza, J. R., Caicedo, A., & D’Abrusco, R. 2024, MNRAS, 528, 4852
- Richards et al. (2002) Richards, G. T., Fan, X., Newberg, H. J., et al. 2002, AJ, 123, 2945
- Roster et al. (2024) Roster, W., Salvato, M., Krippendorf, S., et al. 2024, A&A, 692, A260
- Rozemberczki et al. (2022) Rozemberczki, B., Watson, L., Bayer, P., et al. 2022, arXiv e-prints, arXiv:2202.05594
- Salpeter (1964) Salpeter, E. E. 1964, ApJ, 140, 796
- Salvato et al. (2018) Salvato, M., Buchner, J., Budavári, T., et al. 2018, MNRAS, 473, 4937
- Salvato et al. (2008) Salvato, M., Hasinger, G., Ilbert, O., et al. 2008, ApJ, 690, 1250
- Salvato et al. (2011) Salvato, M., Ilbert, O., Hasinger, G., et al. 2011, ApJ, 742, 61
- Salvato et al. (2019) Salvato, M., Ilbert, O., & Hoyle, B. 2019, Nature Astronomy, 3, 212
- Salvato et al. (2022) Salvato, M., Wolf, J., Dwelly, T., et al. 2022, A&A, 661, A3
- Sani et al. (2011) Sani, E., Marconi, A., Hunt, L. K., & Risaliti, G. 2011, MNRAS, 413, 1479
- Saxena et al. (2024) Saxena, A., Salvato, M., Roster, W., et al. 2024, A&A, 690, A365
- Shakura & Sunyaev (1973) Shakura, N. I. & Sunyaev, R. A. 1973, A&A, 24, 337
- Shankar et al. (2016) Shankar, F., Bernardi, M., Sheth, R. K., et al. 2016, MNRAS, 460, 3119
- Silva et al. (2016) Silva, D. R., Blum, R. D., Allen, L., et al. 2016, in AAS, Vol. 228, AAS Meeting Abstracts #228, 317.02
- Smolčić et al. (2017) Smolčić, V., Novak, M., Delvecchio, I., et al. 2017, A&A, 602, A6
- Stern et al. (2012) Stern, D., Assef, R. J., Benford, D. J., et al. 2012, ApJ, 753, 30
- Suh et al. (2020) Suh, H., Civano, F., Trakhtenbrot, B., et al. 2020, ApJ, 889, 32
- Sutherland & Saunders (1992) Sutherland, W. & Saunders, W. 1992, MNRAS, 259, 413
- Temple et al. (2021) Temple, M. J., Hewett, P. C., & Banerji, M. 2021, MNRAS, 508, 737
- Traulsen et al. (2020) Traulsen, I., Schwope, A. D., Lamer, G., et al. 2020, A&A, 641, A137
- Urry & Padovani (1995) Urry, C. M. & Padovani, P. 1995, PASP, 107, 803
- Vallenari et al. (2023) Vallenari, A., Brown, A. G. A., Prusti, T., et al. 2023, A&A, 674, A1
- Veilleux & Osterbrock (1987) Veilleux, S. & Osterbrock, D. E. 1987, ApJS, 63, 295
- Webb et al. (2020) Webb, N. A., Coriat, M., Traulsen, I., et al. 2020, A&A, 641, A136
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
- Xue et al. (2011a) Xue, Y. Q., Brandt, W. N., Luo, B., et al. 2011a, in ASP, Vol. 439, The Galactic Center: a Window to the Nuclear Environment of Disk Galaxies, ed. M. R. Morris, Q. D. Wang, & F. Yuan, 478
- Xue et al. (2011b) Xue, Y. Q., Luo, B., Brandt, W. N., et al. 2011b, ApJS, 195, 10
- Zeraatgari et al. (2024) Zeraatgari, F. Z., Hafezianzadeh, F., Zhang, Y., et al. 2024, MNRAS, 527, 4677
- Zou et al. (2017) Zou, H., Zhang, T., Zhou, Z., et al. 2017, ApJ, 153
Appendix A RF parameters
Understanding which features are most relevant to our model estimates is crucial for predictive accuracy. An ML model operating as a black box may produce predictions, but the lack of transparency raises questions about the reliability of the outcomes. This not only hinders immediate interpretability but also complicates post-analyses. The ML branch investigating the role each feature plays in the model’s predictions is commonly referred to as ’feature importance’. To this end, we compute SHapley Additive exPlanations (SHAP; Lundberg & Lee 2017; Rozemberczki et al. 2022; Chen et al. 2022) values, based on a game-theoretic approach, by calculating the partial relevance for each feature. Based on a conditional expectation function, features are assigned importance values given their impact on test sample predictions. For a total achievable value , SHAP values are computed via
Given a given feature , the summation is taken over all subsets of the feature set , which one can construct after excluding . The size of is given by while and refer to the value achieved by per subset before and after joins , respectively. This ensures that the contribution of each feature to a prediction collectively adds up to the value of the prediction as
However, achieving optimal accuracy in large modern data sets often involves complex models such as ensembles or deep-learning frameworks. Since it would be unfeasible to compute contributions across all possibilities of the feature space, we approximate SHAP values through Deep SHAP (Rozemberczki et al. 2022; Chen et al. 2022), which efficiently combines values computed for smaller network components into comprehensive values for the entire network, avoiding the need for heuristic choices in linearising components. The sign of SHAP values denotes whether a feature contributes to an increase or decrease in the predictive output, while the magnitude of the SHAP value quantifies the extent of the feature’s influence on the prediction. Features with SHAP values close to zero for individual sources can be considered negligible or non-influential in this scenario. As a consequence, a feature can be seen as most influential if it exerts the most influence on the output across all predictions (Dey et al. 2022).
Appendix B Column description of released CTP catalogue
1.
Xray_ID: X-ray source identifier (unique)
2.
Xray_RA: J2000 Right Ascension of the X-ray source in the parent catalogue (deg)
3.
Xray_DEC: J2000 Declination of the X-ray source in the parent catalogue (deg)
4.
Xray_CAT: X-ray source catalogue
5.
Xray_pos_err: Positional uncertainty of X-ray source (arcsec)
6.
Fx: Mean 0.5 – 2 keV X-ray flux from all detections of the X-ray source ()
7.
EFx: Error on the mean 0.5 – 2 keV X-ray flux from all detections of the X-ray source ()
8.
S_Nx: Signal-to-noise ratio for Fx/EFx
9.
Fx_hard: Mean hard (4XMM DR14: 2 – 4.5 keV, CSC 2.0: 2 – 7 keV and eROSITA DR1: 2–5 keV) X-ray flux from all detections of the X-ray source ()
10.
EUCLID_ID: EUCLID source identifier (unique)
11.
EUCLID_RA: J2000 Right Ascension of the counterpart in Q1 catalogue (deg)
12.
EUCLID_DEC: J2000 Declination of the counterpart in Euclid catalogue (deg)
13.
EUCLID_pos_err: Positional uncertainty of Euclid source (arcsec)
14.
Xray_EUCLID_ID: Concatenation of X-ray_ID and EUCLID_ID (unique)
15.
EUCLID_Xray_proba: Probability to be an X-ray emitter based on the RF
16.
Separation_EUCLID_Xray: Distance between Euclid and X-ray source (arcsec)
17.
dist_bayesfactor: Logarithm of ratio between prior and posterior, from separation, positional error, and number density (see Appendix in Salvato et al. 2018)
18.
dist_post: Distance posterior probability comparing this association vs. no association (see Appendix in Salvato et al. 2018)
19.
bias_EUCLID_Xray_proba: Probability weighting introduced by the Euclid-based prior (1 indicates no change)
20.
p_single: Same as dist_post, but weighted by the prior (see Appendix in Salvato et al. 2018)
21.
p_any: Probability that there is a counterpart in LS10 for each X-ray entry (see Appendix in Salvato et al. 2018)
22.
p_i: Relative probability of the X-ray/Euclid match (see Appendix in Salvato et al. 2018)
23.
match_flag: 1 for the most probable match; 2 for almost as good solutions (p_i/p_ibest ¿ 0.5)
24.
EDF: Euclid deep field
25.
LS10_FULLID: Unique identifier of the LS10 source matched to the Euclid CTP (concatenation of RELEASE, BRICKID, and OBJID)
26.
LS10_RA: J2000 Right Ascension of the LS10 source (deg)
27.
LS10_DEC: J2000 Declination of the LS10 source (deg)
28.
Gal_proba: Probability of being a Galactic rather than extragalactic source based on RF
29.
ra_spec: J2000 Right Ascension of the spec- match to the Euclid CTP (deg)
30.
dec_spec: J2000 Declination of the spec- match to the Euclid CTP (deg)
31.
z_spec: Spectroscopic redshift
32.
z_err: Error on the spectroscopic redshift
33.
Cat_spec: Parent catalogue of the spec- match
34.
phz: PICZL photometric redshift (see Roster et al. 2024)
35.
phz_1sl: Photo- lower 1 sigma error
36.
phz_1su: Photo- upper 1 sigma error
37.
phz_3sl: Photo- lower 3 sigma error
38.
phz_3su: Photo- upper 3 sigma error
39.
z_final: Spec- when available; otherwise, photo-
40.
Lx: X-ray luminosity, assuming FlatLambdaCDM from astropy.cosmology

### v2

@@
 point-like X-ray selected sources
-To better understand the role of active galactic nuclei (AGN) in galaxy evolution, it is crucial to achieve a complete and pure AGN census.
-X-ray surveys are key to this, but identifying their counterparts (CTPs) at other wavelengths remains challenging due to their larger positional uncertainties and limited availability of deeper, uniform ancillary data.
-Euclid is revolutionising this effort, offering extensive coverage of nearly the entire extragalactic sky, particularly in the near-infrared bands, where AGN are more easily detected.
-With the first Euclid Quick Data Release (Q1), we tested and validated methods for identifying, classifying, and determining the redshifts of Euclid CTPs to known point-like sources from major X-ray surveys, including XMM-Newton, Chandra, and eROSITA.
-Using Bayesian statistics, combined with machine learning (ML), as incorporated in the algorithm NWAY, we identify the CTPs to 11 286 X-ray sources from the three X-ray telescopes.
+To better understand the role of active galactic nuclei (AGN) in galaxy evolution, it is crucial to work with a complete and pure AGN sample.
+X-ray surveys are key to doing so, but their larger positional uncertainties complicate counterpart (CTP) association, further compounded by the limited availability of deep, uniform multi-wavelength ancillary data.
+Euclid is revolutionising this identification effort, offering extensive coverage of nearly the entire extragalactic sky, particularly in the near-infrared bands, where AGN are more easily detected.
+Using the first Euclid Quick Data Release (Q1), we validated the methods for identifying and classifying Euclid CTPs of known point-like sources from major X-ray surveys, including XMM-Newton, Chandra, and eROSITA.
+Using Bayesian statistics, combined with machine learning (ML), as incorporated in the algorithm NWAY, we identified the CTPs of 11 286 X-ray sources from the three X-ray telescopes.
 For the large majority of 10 194 sources, the association is unique, with the remaining 10% of multi-CTP cases equally split between XMM-Newton and eROSITA.
 Six percent of the Euclid CTPs are detected in more than one X-ray survey.
-ML is then used to distinguish between Galactic (8%) and extragalactic (92%) sources.
-We computed photo-s using deep learning for the 8617 sources detected in the 10th data release of the DESI Legacy Survey, reaching an accuracy and a fraction of outliers of about 5%.
-Based on their X-ray luminosities, over 99% of CTPs identified as extragalactic are classified as AGN, most of which appear unobscured according to their hardness ratios.
-With this paper, we release our catalogue, which includes identifiers, basic X-ray properties, the reliability of the associations, and additional property extensions such as Galactic/extragalactic classifications and photometric/spectroscopic redshifts.
-We also provide probabilities for sub-selecting the sample based on purity and completeness, allowing users to tailor the sample according to their specific needs.
+We then used ML to distinguish between Galactic (8%) and extragalactic (92%) sources.
+We computed photo-s using deep learning for the 9259 sources detected in the tenth data release of the DESI Legacy Survey, reaching an accuracy and a fraction of outliers of roughly 5%.
+Based on their X-ray luminosities, all CTPs identified as extragalactic are classified as AGN, most of which appear as type I AGN according to their hardness ratios.
+With this paper, we release our catalogue, which includes identifiers, basic X-ray properties, the reliability of the associations, and additional property extensions, such as Galactic- or extragalactic classifications and photometric/spectroscopic redshifts.
+We also provide probabilities for sub-selecting the sample based on purity and completeness, in order to allow users to tailor the sample according to their specific needs.
 Key Words.:
 Methods: statistical, data analysis; Surveys; Catalogues; Galaxies: active; X-rays: galaxies1 Introduction
-Active galactic nuclei (AGN111Used interchangeably in both singular and plural cases), which rank among the most energetic phenomena in the Universe, play a pivotal role in shaping the evolutionary trajectories of galaxies throughout cosmic time.
+Active galactic nuclei (AGN111We use this term interchangeably to indicate both the singular and plural cases.), which rank among the most energetic phenomena in the Universe, play a pivotal role in shaping the evolutionary trajectories of galaxies throughout cosmic time.
 Their activity is fundamentally driven by the accretion of matter onto supermassive black holes (SMBHs) from their surrounding environment (e.g., Salpeter 1964; Lynden-Bell 1969; Pringle & Rees 1972; Shakura & Sunyaev 1973; Lynden-Bell & Pringle 1974; Kormendy & Ho 2013).
@@
 2017).
-While this is particularly true for quasi-stellar objects (QSOs), it is important to note that the range of activity, emission, and interaction of AGN spans a broad spectrum, with different AGN exhibiting varying levels of energy outputs (e.g., Peterson 1997; Pierce et al.
+While this is particularly true for quasi-stellar objects (QSOs), it is important to note that the range of activity, emission, and interaction of AGN span a broad spectrum, with different AGN exhibiting varying levels of energy outputs (e.g., Peterson 1997; Pierce et al.
 2010; Pović et al.
@@
 These close connections are supported by fundamental scaling relations that link the SMBH mass to various galactic properties.
-Notable among these are the – relation, which ties BH mass to the velocity dispersion of stars in the galaxy’s bulge (Gebhardt et al.
-2000; Ferrarese & Merritt 2000), and the – relation, which connects it to the stellar mass of the host galaxy (Magorrian et al.
+Notable among these are the - relation, which ties black hole (BH) mass to the velocity dispersion of stars in the galaxy’s bulge (Gebhardt et al.
+2000; Ferrarese & Merritt 2000), and the - relation, which connects it to the stellar mass of the host galaxy (Magorrian et al.
 1998; Häring & Rix 2004; Gültekin et al.
@@
 2016; Suh et al.
-2020), pointing to an intricate co-evolution of the systems (Kormendy & Ho 2013; Heckman & Best 2014; Madau & Dickinson 2014; Euclid Collaboration: Bisigello et al.
+2020), thus pointing to an intricate co-evolution of the systems (Kormendy & Ho 2013; Heckman & Best 2014; Madau & Dickinson 2014; Euclid Collaboration: Bisigello et al.
 2024).
@@
 2017; Husemann & Harrison 2018; Harrison & Ramos Almeida 2024).
-Despite significant advancements over the years, the current AGN census remains incomplete due to inherent biases introduced to individual samples by various multi-wavelength selection methods, each favouring specific properties of the population (Messias et al.
+Despite significant advancements over the years, the current AGN census remains incomplete due to inherent biases introduced to individual samples by various multi-wavelength selection methods, as each favours specific properties of the population (Messias et al.
 2014; Padovani et al.
@@
 2024).
-Comprehensive wavelength coverage is essential for identifying diverse AGN types across all redshifts, with each method capitalising on the dominant physical processes within its waveband to uncover unique facets of AGN activity, where the AGN dominates the spectral energy distribution (SED) of its host galaxy (Padovani et al.
+Comprehensive wavelength coverage is vital for identifying diverse AGN across redshift, as each method that contributes to this coverage capitalises on the dominant physical processes of AGN activity, dominating the spectral energy distribution (SED) of its host galaxy (Padovani et al.
 2017).
-For instance, radio observations are effective in identifying jet-dominated AGN (De Breuck et al.
+For instance, radio observations are effective at identifying jet-dominated AGN (De Breuck et al.
 2002; Smolčić et al.
@@
 2012; Assef et al.
-2013), which penetrate dense dust clouds effectively (Hickox & Alexander 2018).
-Notably, the significantly higher contrast between the intrinsic brightness of accreting SMBHs and their host galaxies in X-rays, as opposed to other wavelengths, makes this selection one of the most uncontaminated, providing pure AGN samples of larger diversity (Xue et al.
+2013), which can effectively penetrate dense dust clouds (Hickox & Alexander 2018).
+Notably, the significantly higher contrast between the intrinsic brightness of accreting SMBHs and their host galaxies in X-rays, as opposed to other wavelengths, makes this selection one of the most uncontaminated, and thus it provides pure AGN samples of a larger diversity (Xue et al.
 2011a; Donley et al.
@@
 To combine different observational perspectives of AGN, these X-ray samples must then be paired with ancillary multi-wavelength data.
-In the past, identifying multi-wavelength counterparts (CTPs) for X-ray selected sources was a significant hurdle in AGN characterisation and redshift estimation (Salvato et al.
+In the past, identifying multi-wavelength counterparts (CTPs) for X-ray-selected sources was a significant hurdle in AGN characterisation and redshift estimation (Salvato et al.
 2018).
 While combining data from multiple wavelengths is crucial for constructing a more complete picture of AGN, the process of cross-identifying sources remains complicated.
-Due to the significant positional uncertainties in X-ray surveys, confidently linking most X-ray sources to a single, definitive CTP in e.g., optical, infrared, or radio surveys is challenging, except for the brightest cases.
+Due to the significant positional uncertainties in X-ray surveys, confidently linking most X-ray sources to a single, definitive CTP in, for example, optical, infrared, or radio surveys, is challenging, except in the brightest cases.
 These uncertainties, along with the uneven spatial coverage of ancillary data, further complicate the task of reliably pinpointing AGN across different surveys.
@@
 2024).
-While challenging to this date, the European Space Agency’s Euclid satellite will help overcome this limitation due to its uniform depth and wide coverage (Laureijs et al.
+The European Space Agency’s Euclid satellite will also help overcome this limitation due to its uniform depth and wide coverage (Laureijs et al.
 2011; Euclid Collaboration: Mellier et al.
-2024).
-At the same time, Euclid’s depth results in very high source densities, increasing the risk of unforced misidentification and chance alignments.
+2025).
+At the same time, Euclid’s depth results in very high source densities, thus increasing the risk of unforced misidentification and chance alignments.
 Over its six-year mission lifetime, Euclid will cover a total of 14 000 deg2 in the Euclid Wide Survey (EWS, Euclid Collaboration: Scaramella et al.
-2022), expected to detect billions of sources.
-In addition to its cosmological objectives, Euclid will play a crucial role in the detection and characterization of AGN, expecting to detect at least 10 million of them (Euclid Collaboration: Bisigello et al.
+2022), and it is expected to detect billions of sources.
+In addition to its cosmological objectives, Euclid will play a crucial role in the detection and characterisation of AGN, as at least 10 million AGN are anticipated to be detected (Euclid Collaboration: Bisigello et al.
 2024; Euclid Collaboration: Lusso et al.
@@
 2020; Euclid Collaboration: Cropper et al.
-2024), covering wavelengths from 0.53 to 0.90 () and the Near-Infrared Spectrometer and Photometer (NISP, Maciaszek et al.
+2025), covering wavelengths from 0.53 to 0.90 (), and the Near-Infrared Spectrometer and Photometer (NISP, Maciaszek et al.
 2022; Euclid Collaboration: Jahnke et al.
-2024) operating in the near-infrared, with photometric measurements in three bands (, , and ), covering in total the wavelength range from 0.95 to 2.02 , while also including a slitless spectrograph for the detection of emission lines.
-With the first Euclid Quick Data Release (Q1) we can now begin building and testing the necessary machinery for identifying and characterising AGN previously selected in the X-rays, where we aim to leverage Euclid’s depth, resolution and wavelength coverage to provide more secure associations.
-This initial testing will form the foundation for refining our methods and ensuring accurate AGN identification as more data becomes available over the next years.
+2025).
+Operating in the near-infrared and with photometric measurements in three bands (, , and ), Euclid covers in total the wavelength range from 0.95 to 2.02 while also including a slitless spectrograph for the detection of emission lines.
+With the first Euclid Quick Data Release (Q1), we can now begin building and testing the necessary machinery for identifying and characterising AGN previously selected in the X-rays.
+We aim to leverage Euclid’s depth, resolution, and wavelength coverage to provide more secure associations of these AGN.
+Moreover, this initial testing will form the foundation for refining our methods and ensuring accurate AGN identification as more data become available over the next years.
 In this paper, we address the task of reliably associating X-ray sources with their Euclid CTPs and investigate their respective NIR properties as a function of X-ray flux, among other properties.
-The paper is structured as follows: Sect.
-2 gives an overview of the Q1 data.
-Section 3 introduces the X-ray source catalogues used in this work, while Sect.
-4 outlines the approach to identifying CTPs.
-Section 5 describes how using a sample of sources clearly characterised with the 10th Data Release of the Legacy Survey (Dey et al.
-2019), we assign a Galactic/extragalactic nature to our sources using Euclid photometry.
+The paper is structured as follows: In Sect.
+2 we give an overview of the Q1 data.
+In Sect.
+3 we introduce the X-ray source catalogues used in this work, while Sect.
+4 provides an outline of the approach to identifying CTPs.
+In Sect.
+5, using a sample of sources clearly characterised with the tenth Data Release of the Legacy Survey (Dey et al.
+2019), we describe how we assign a Galactic- or extragalactic nature to our sources using Euclid photometry.
 In Sect.
 6, we compute photometric redshifts for all CTPs.
-Section 7 displays CTP sample properties and Sect.
-8 explains the release of the catalogues and how to use them.
-A summary and discussion are given in the final Sect.
+Section 7 displays CTP sample properties, and in Sect.
+8 we explain the release of the catalogues and how to use them.
+Finally, a summary and discussion are given in Sect.
 9.
-In this paper, unless stated otherwise, we express magnitudes in the AB system (Oke & Gunn 1983) and adopt a flat CDM cosmology with km s-1 Mpc-1, and , to facilitate direct comparison with similar works from literature.
+In this paper, unless stated otherwise, we express magnitudes in the AB system (Oke & Gunn 1983) and adopt a flat CDM cosmology with km s-1 Mpc-1, , and to facilitate direct comparison with similar works from literature.
 2 Euclid Q1 fields
 Euclid Q1 marks the initial data release with a footprint area of 63.1 deg2 (Euclid Collaboration: Mellier et al.
-2024).
+2025).
 It consists of observations making up the Euclid Deep Survey (EDS), selected for their considerable multi-wavelength coverage.
 However, the data is currently at the depth of the EWS, with a 5 depth of (AB)24.5 for point sources.
-The EDS targets three important Euclid Deep Fields (EDFs, see Fig.
+The EDS targets three important Euclid Deep Fields (EDFs; see Fig.
 1): the EDF Fornax (EDF-F), EDF South (EDF-S), and the EDF North (EDF-N).
@@
 2020) and Chandra Source Catalogue 2 (CSC 2.0, Evans et al.
-2024) feature outstanding sensitivity for X-ray selected sources but cover relatively small regions of the sky.
-On the other hand, wide-area surveys are crucial for complementing pointed observations and uncover those AGN missed (e.g., Ananna et al.
+2024) feature outstanding sensitivity for X-ray-selected sources but cover relatively small regions of the sky.
+On the other hand, wide-area surveys are crucial for complementing pointed observations and uncovering those AGN missed (e.g., Ananna et al.
 2017; Fotopoulou et al.
@@
 As part of the ongoing eROSITA All-Sky Survey (eRASS), these data provide homogeneous, soft X-ray coverage of the entire extragalactic sky.
-While eROSITA’s positional accuracy is somewhat pourer than that of XMM-Newton and Chandra, its unparalleled breadth allows us to include a significantly larger sample of sources, enhancing the statistical power of our analysis and ensuring a more comprehensive AGN census (see Fig.
+While eROSITA’s positional accuracy is somewhat poorer than that of XMM-Newton and Chandra, its unparalleled breadth allows us to include a significantly larger sample of sources, enhancing the statistical power of our analysis and ensuring a more comprehensive AGN census (see Fig.
 1 and Footnote 2).
 In the following subsections, we describe the selection criteria applied to the catalogues from each X-ray survey.
-X-ray catalogue | Number of sources | AND point-like | AND POSERR ¡ limit | EDF-F | EDF-N | EDF-S |
-(1) | (2) | (3) | (4) | (5) | (6) | (7) |
-4XMM DR14 | 692 109 | 612 521 | 612 504 | 4523 | 445 | 273 |
-CSC 2.0 | 317 167 | 296 473 | 296 162 | 1016 | 506 | 147 |
-eROSITA DR1 main | 898 812 | 879 153 | 878 984 | 1012 | 0 | 3369 |
+| X-ray catalogue | MOC | Number of sources | AND point-like | AND POSERR ¡ limit | EDF-F | EDF-N | EDF-S |
+| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
+| Unit | deg2 | ||||||
+| 4XMM DR14 | 6.18 | 692 109 | 612 521 | 612 504 | 4523 | 445 | 273 |
+| CSC 2.0 | 1.31 | 317 167 | 296 473 | 296 162 | 1016 | 506 | 147 |
+| eROSITA DR1 main | 43.14 | 898 812 | 879 153 | 878 984 | 1012 | 0 | 3369 |
 3.1 XMM-Newton
 The XMM-Newton space-telescope provides some of the deepest and most detailed X-ray imaging available, with a spatial resolution of approximately 6″ and positional uncertainties typically below 1–2″.
-The data used in this study include both targeted and serendipitous observations, selected from the 4XMM-DR14 catalogue333catalogue and data model description available at https://cdsarc.cds.unistra.fr/viz-bin/cat/IX/69, which offers multi-band flux measurements and quality flags.
+The data used in this study includes both targeted and serendipitous observations, selected from the 4XMM DR14 catalogue333catalogue and data model description available at https://cdsarc.cds.unistra.fr/viz-bin/cat/IX/69, the most recent of the XMM releases444We note that dedicated XMM observations targeting the Fornax field are ongoing.
+These additional data are not included in this work but may be used to refine and extend the CTPs in future..
+The catalogue offers multi-band flux measurements and quality flags.
 To prepare the sample, point-like sources were filtered based on their extent likelihood and positional uncertainties: SC_EXT_ML 10, SC_EXTENT 10 and SC_POSERR 10″, where the threshold for the positional uncertainty is motivated by the respective distribution as a function of flux as shown in Fig.
@@
 This effectively reduces the computational load for cross-matching with Euclid.
-To be consistent with other X-ray surveys, we compute the X-ray flux for the 0.5 – 2 keV band as
+To be consistent with other X-ray surveys, we computed the X-ray flux, , for the 0.5–2 keV band as
 and its uncertainty as
-by combining the 0.5 – 1 keV, and 1 – 2 keV band fluxes and errors.
+by combining the 0.5–1 keV and 1–2 keV band fluxes and errors.
 3.2 Chandra
-The Chandra X-ray space-telescope provides sub-arcsecond angular resolution, which allows precise localisation of sources and makes it particularly effective in crowded or complex fields.
-Our sample incorporates sources from the CSC 2.0 catalogue444https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=IX/57&-to=3.
-To subsample only point-like sources, we select sources with the extended flag, , set to zero.
-The 0.5 – 2 keV X-ray flux and its uncertainty are computed as
+The Chandra X-ray space telescope provides sub-arcsecond angular resolution, which allows for precise localisation of sources and makes it particularly effective in crowded or complex fields.
+Our sample incorporates sources from the CSC 2.0 catalogue555https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=IX/57&-to=3.
+To subsample only point-like sources, we selected sources with the extended flag, , set to zero.
+The 0.5–2 keV X-ray flux and its uncertainty were computed as
 with Flux and Flux referring to the default CSC 2.0 soft and medium bands, respectively, while (B) and (b) represent the upper and lower 1 error margins.
-The positional uncertainties are determined using the 2 errors along the semi-major (0) and semi-minor (1) axes of the error ellipse provided in the source catalogue, as
-and downsample the CSC 2.0 catalogue such that C_POSERR 10″, to be consistent with 4XMM DR14 (see Fig. 2).
+The positional uncertainties were determined using the 2 errors along the semi-major (0) and semi-minor (1) axes of the error ellipse provided in the source catalogue as
+and we downsampled the CSC 2.0 catalogue such that C_POSERR 10″to be consistent with 4XMM DR14 (see Fig. 2).
 3.3 eROSITA
-The eROSITA-DE DR1 main catalogue555https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/Catalogues_dr1/ offers comprehensive coverage of the western galactic hemisphere.
+The eROSITA-DE DR1 main catalogue666https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/Catalogues_dr1/ offers comprehensive coverage of the western galactic hemisphere.
 While it does not include the EDF-N region (see Fig.
@@
 Until recently, the identification of CTPs was primarily performed using maximum likelihood matching (e.g., Sutherland & Saunders 1992).
-This approach considers the separation between sources in the primary catalogue (e.g., X-ray) and the secondary multiwavelength catalogue, the positional uncertainties of the primary sources, and the magnitude distribution of sources in the ancillary catalogues within a certain radius from the primary sources.
+This approach considers the separation between sources in the primary catalogue (e.g.
+X-ray) and the secondary multiwavelength catalogue, the positional uncertainties of the primary sources, and the magnitude distribution of sources in the ancillary catalogues within a certain radius from the primary sources.
 While this method has been applied successfully in cases with small positional errors and bright sources (see e.g., Naylor et al.
@@
 4.1 NWAY
-To address these challenges, bayesian statistic needs to be invoked (Budavári & Szalay 2008).
+To address these challenges, Bayesian statistics needs to be invoked (Budavári & Szalay 2008).
 With this in mind, algorithms such as NWAY (Salvato et al.
@@
 It refines its matching procedure by being able to supplement one or more features, such as colours, magnitudes, and S/N, into likelihood ratios used for CTP identification.
-These priors, based on (anti-) correlations between source properties of the target and field populations, allow NWAY to improve the likelihood of finding true matches, boosting the accuracy significantly over methods that rely solely on positional information.
+These priors based on (anti-) correlations between source properties of the target and field populations allow NWAY to improve the likelihood of finding true matches, boosting the accuracy significantly over methods that rely solely on positional information.
 4.1.1 The integration of priors in NWAY
 Colour and magnitude priors are posteriors that encode knowledge from data or probabilistic distributions that describe the expected properties of true CTPs for the sources in the primary catalogue.
-For example, an AGN selected in X-rays is likely to have distinct colours (e.g., redder MIR colours due to dust) and magnitudes compared to stars or inactive galaxies.
+For example, an AGN selected in X-rays is likely to have distinct colours (e.g.
+redder MIR colours due to dust) and magnitudes compared to stars or inactive galaxies.
 The difference between the methods resides then in the adoption of specific features able to distinguish an X-ray emitter (regardless of its Galactic or extragalactic nature) from a random source in the field (Salvato et al.
@@
 2022), Salvato et al.
-(2022) defined a prior that utilised not only magnitudes and colours from the Legacy Survey Data Release 10 (Dey et al.
+(2022) defined a prior that utilised not only magnitudes and colours from the Legacy Survey Data Release 10 (LS10, Dey et al.
 2019), but also S/N in all bands (griz, W1, W2, W3, and W4), as well as properties, such as proper motion, from Gaia DR3 (Vallenari et al.
 2023) when available.
-For further details on the formalism, we refer to the NWAY documentation777https://github.com/JohannesBuchner/NWAY.
+For further details on the formalism, we refer to the NWAY documentation.888https://github.com/JohannesBuchner/NWAY
 4.1.2 Q1 prior
@@
 2011).
-We start by utilising a sample of secure CTPs to X-ray sources from 4XMM DR11 (Traulsen et al.
-2020) and CSC 2.0 for training the RF as presented by Salvato et al.
-(2022).
+We start by utilising the training sample presented in Salvato et al., in prep., comprising secure CTPs to X-ray sources from 4XMM DR11 (Webb et al.
+2020) and CSC 2.0, for training the RF model, as presented by (Salvato et al.
+2022).
 NWAY is then applied to identify CTPs for these X-ray sources in the EDF catalogues.
@@
 The sky coverage of each X-ray sample per EDF is calculated to account for overlapping search windows.
-Specifically, while the coverage for the X-ray sample is given by its multi-order coverage map (MOC) or pointing, the EDF coverage is derived as
+Specifically, while the coverage for the X-ray sample is given by its multi-order coverage map (MOC) or pointing, the EDF coverage was derived as
 where corresponds to the number of X-ray sources and is the overlapping area of adjacent search windows.
@@
 This field population serves as a representation of non-X-ray sources, providing a comparison set for training the RF.
-To label the data sets, X-ray CTPs are assigned a target class of ’1’, while field objects are labelled as ‘0‘.
-The Euclid-only training features utilised for the RF classifier are outlined in Footnote 6.
+To label the data sets, X-ray CTPs are assigned a target class of ‘1’, while field objects are labelled as ‘0‘.
+The Euclid-only training features utilised for the RF classifier are outlined in Footnote 7.
 At present, photometry and colours from different apertures are not included as features, as the aperture sizes vary between sources despite sharing the same designation.
 Future data releases may standardise aperture sizes across all objects, also enhancing the RF classifier’s ability to leverage the relationship between light profiles and X-ray emission.
-To evaluate model performance, 15% of the training data set (6000 field and 1730 target sources) is randomly set aside as a test set.
+To evaluate model performance, 15% of the training dataset (6000 fields and 1730 target sources) is randomly set aside as a test set.
 The RF model is configured with 200 decision trees, with splits allowed if at least 25 samples remain in a branch.
-Per decision tree construction, 10 photometric features are considered and bootstrap sampling is applied to the training set for building the ensemble of trees.
+Per decision tree construction, 10 photometric features are considered, and bootstrap sampling is applied to the training set for building the ensemble of trees.
 Since the data set is highly imbalanced, with field objects vastly outnumbering CTPs of X-ray sources, a weighting scheme is implemented to automatically adjust the contribution of each class during training (Salvato et al.
@@
 3, assessed by acquiring these values for the test set (see e.g., Fotopoulou & Paltani 2018).
-We define the following measures of quality for the classifier:
+We defined the following measures of quality for the classifier:
 accuracy (TPTN)/(TPTNFPFN) ;
@@
 fall-out FP/(TNFP) .
-The trained model achieves strong performance, with high accuracy of 89% and recall of 91%, indicating that most real X-ray emitters are correctly identified.
+The trained model achieves strong performance, with a high accuracy of 89% and a recall of 91%, indicating that most real X-ray emitters are correctly identified.
 The contamination rate remains low, with a fractional leakage of 11%.
@@
 | in LS10 | S/Nr ¿= 3 | ||||||
-4XMM DR14 | 5241 | 5885 | 5239 | 3774 | 3711 | 5486 | 1122 |
-CSC 2.0 | 1669 | 1712 | 1666 | 684 | 669 | 1625 | 442 |
-eROSITA DR1 | 4381 | 5048 | 4381 | 4159 | 4141 | 4375 | 477 |
-Total: | 11 291 | 12 645 | 11 286 | 8617 | 8521 | 11 486 | 2044 |
+| 4XMM DR14 | 5241 | 5885 | 5239 | 4107 | 4033 | 5486 | 953 |
+| CSC 2.0 | 1669 | 1712 | 1666 | 993 | 958 | 1625 | 502 |
+| eROSITA DR1 | 4381 | 5048 | 4381 | 4159 | 4141 | 4375 | 321 |
+| Total: | 11 291 | 12 645 | 11 286 | 9259 | 9132 | 11 486 | 1776 |
 4.2 Adding the X-ray emitter prior to NWAY
-The a-priori probability of an Euclid source to be an X-ray emitter, , is now incorporated as a prior into the NWAY framework to enhance our CTP identification process.
+The a priori probability of a Euclid source being an X-ray emitter, , is now incorporated as a prior into the NWAY framework to enhance our CTP identification process.
 To visualise the positive impact of the inclusion of this prior, we plot the distribution of p_any both before and after the inclusion of the X-ray emitter prior (see Fig.
@@
 Notably, the inclusion of the prior also changes the respective CTP in about 30% of the cases, indicating that for those sources that are indeed not the CTP after adding the prior, the p_any value has been decreased.
-In total, we associate 12 645 CTP candidates.
-Among these, 11 286 (about ) sources have , indicating that they have the highest probability p_i to be considered reliable CTPs.
+In total, we associated 12 645 CTP candidates.
+Among these, 11 286 sources (about ) have , indicating that they have the highest probability p_i to be considered reliable CTPs.
 The remaining candidates, with , represent secondary but plausible matches that are not ruled out entirely.
-We find 1092 X-ray sources with multiple CTPs, the vast majority of which is made up by XMM-Newton and eROSITA with only Chandra providing the required positional accuracy for Euclid to have few sources with multiple plausible CTPs.
-While most of the multi-CTP cases are made up of two candidates, a few cases with CTPs are introduced, predominantly by eROSITA, given its larger positional uncertainty.
+We found 1092 X-ray sources with multiple CTPs, the vast majority of which are made up by XMM-Newton and eROSITA, with only Chandra providing the required positional accuracy for Euclid to have a few sources with multiple plausible CTPs.
+While most of the multi-CTP cases are made up of two candidates, a few cases with more than two CTPs are introduced, predominantly by eROSITA, given its larger positional uncertainty.
 Notably, only five X-ray sources within the Q1 footprint (see Footnote 2) lack an association entirely, possibly due to boundary effects where the true CTP falls outside the region covered by the EDFs.
-With this catalogue, several scenarios emerge in terms of the relationships between CTPs and X-ray sources.
+With this catalogue, several scenarios emerged in terms of the relationships between CTPs and X-ray sources:
 1.
-Two X-ray entries from differing surveys are assigned to a single CTP.
+Two X-ray entries from different surveys are assigned to a single CTP.
 This scenario suggests but does not confirm the possibility of a single physical source being detected by two different X-ray surveys.
@@
 3.
-Cases where a single Euclid ID is linked to multiple entries from e.g., two or even all three X-ray surveys.
+Cases where a single Euclid ID is linked to multiple entries from two or even all three X-ray surveys.
 4.
@@
 At this stage, gaps in coverage further limit the significance of overlap checks.
-Keeping this in mind we find 674 pairs which either correspond to case one or two, as well as 16 triplets referring to case three as shown in Fig.
+Keeping this in mind, we find 674 pairs which either correspond to case one or two, as well as 16 triplets referring to case three, as shown in Fig.
 5.
 We deliberately choose not to quantify case four, as there is currently no definitive way to resolve it with certainty.
-A breakdown of the CTP sample is given in Footnote 8.
+A breakdown of the CTP sample is given in Footnote 9.
 4.3 Purity and completeness
-To calibrate a reliable p_any cut-off, we perform the NWAY procedure (using the same prior) on a duplicate sample of the X-ray sources, referred to as ’randoms‘.
-These randoms are generated by displacing the positions of the X-ray sources to lie beyond the search radius, ensuring that any matches with multiwavelength CTPs are purely coincidental.
-By comparing the distribution of p_any values for both the original X-ray sample and the randoms, we can assess the impact of the prior by examining the two histograms in Fig.
+To calibrate a reliable p_any cut-off, we performed the NWAY procedure (using the same prior) on a duplicate sample of the X-ray sources, referred to as ‘randoms’.
+These randoms were generated by displacing the positions of the X-ray sources to lie beyond the search radius, ensuring that any matches with multiwavelength CTPs are purely coincidental.
+By comparing the distribution of p_any values for both the original X-ray sample and the randoms, we could assess the impact of the prior by examining the two histograms in Fig.
 4.
-The figure shows how the p_any value for the randoms is usually lower than for the real sample of X-ray sources, indicating that in a random position in the sky, there are no sources that have the same features of a typical X-ray emitters.
+The figure shows how the p_any value for the randoms is usually lower than for the real sample of X-ray sources, indicating that in a random position in the sky, there are no sources that have the same features as a typical X-ray emitter.
 For the few sources for which p_any remains high, either the prior was not representative enough, or, deeper data in the future will potentially reveal a fainter X-ray source.
-To further assess the impact of incorporating into the NWAY procedure, we compare the p_any values as a function of the separation distance between the primary (X-ray) and secondary (multi-wavelength) sources.
+To further assess the impact of incorporating into the NWAY procedure, we compared the p_any values as a function of the separation distance between the primary (X-ray) and secondary (multi-wavelength) sources.
 Figure 6 presents the separation between the 4XMM sources and the associated CTP in EDF-F as a function of p_any.
-The sources with a higher probability of being correct also have a high probability of being X-ray emitters, even at larger separation, the latter being mostly smaller than 10″.
+The sources with a higher probability of being correct also have a high probability of being X-ray emitters, even at larger separations, the latter being mostly smaller than 10″.
 CTPs of little separation and low p_any showcase scenarios of chance alignments where objects appear close but do not show significant .
-Next, we focus on evaluating the purity and completeness of the CTP samples based on how they are sub-selected.
-In this context, we define completeness as the cumulative fraction of sources with p_any higher than a certain value and purity the fraction of sources with p_any higher than the value in the random sample.
+Next, we focused on evaluating the purity and completeness of the CTP samples based on how they are sub-selected.
+In this context, we defined completeness as the cumulative fraction of sources with p_any higher than a certain value and purity as the fraction of sources with p_any higher than the value in the random sample.
 This gives a measure of how often the CTP in the real sample is associated to a CTP only by chance.
-Figure 7 shows both purity and completeness for each EDF and X-ray survey as well as the ’sweet spot’ for the best trade off between these quantities as a function of p_any.
+Figure 7 shows both the purity and completeness for each EDF and X-ray survey as well as the ‘sweet spot’ for the best trade-off between these quantities as a function of p_any.
 Variations in the distributions result from differences in the positional uncertainties of the X-ray surveys and the source densities within both the X-ray data and the corresponding EDF.
-Depending on the scientific goal, a user interested in creating a catalogue that is extremely complete or pure, is hence free to select sources with a p_any beyond a certain threshold.
+Depending on the scientific goal, a user interested in creating a catalogue that is extremely complete or pure is hence free to select sources with a p_any beyond a certain threshold.
 For the rest of the work, we adopt the intersection between the purity and completeness curves as the threshold.
 In all fields, purity and completeness are at least 80%, with the Chandra catalogues unsurprisingly having the most reliable associations, reaching values of about 90%.
-Interestingly, the impact of the higher spatial resolution of Chandra is clear in the EDF-F field, where Fornax, the second richest galaxy cluster of the local Universe is located.
+Interestingly, the impact of the higher spatial resolution of Chandra is clear in the EDF-F field, where Fornax, the second richest galaxy cluster of the local Universe, is located.
 In such a dense environment, it is more challenging to identify the correct CTP given the resolution of 4XMM and eROSITA.
-To conclude this part, we assess the stability in the method by examining the impact of RF hyper-parameters like minimum leaf size and features per tree.
+To conclude this part, we assessed the stability of the method by examining the impact of RF hyperparameters such as the minimum leaf size and features per tree.
 Variations in these parameters mainly shift the intersection of purity and completeness curves along the p_any axis, with smaller leaf sizes enhancing subclass specificity and subtly altering the intersection point.
@@
 The primary distinction is between extragalactic sources (such as galaxies or faint and bright AGN) and galactic sources (including stars and compact objects).
-In the following section, we outline the methodology used for classifying the sources and the validation tests conducted to ensure the accuracy of the classification, the results of which are summarised in Footnote 8.
+In the following section, we outline the methodology used for classifying the sources and the validation tests conducted to ensure the accuracy of the classification, the results of which are summarised in Footnote 9.
 5.1 Galactic and extragalactic sources
@@
 2017).
-We match the 11 286 CTPs to LS10 using a positional tolerance and find an association for 68% of our sources, see Footnote 8.
+We match the 11 286 CTPs to LS10 using a positional tolerance and find an association for 82% of our sources, see Footnote 9.
 Next, we apply a cut of S/N ¿ 3 in the , , , and W1 bands to minimise the relative impact of photometric uncertainties.
@@
 5.2 Probabilistic classifier approach
-To overcome this limitation, we follow the methodology outlined in Sect.
-4.1.1, training an RF classifier using the same Euclid-only features listed in Footnote 6.
-However, this time we label the sample based on the classification derived from the versus plot, setting Galactic objects (396) to ‘1‘ and extragalactic objects (1000) to ‘0‘.
-Using the trained model, we assign a Galactic or extragalactic label to each Euclid CTP, providing a probabilistic classification for robust downstream analyses.
+To overcome this limitation, we followed the methodology outlined in Sect.
+4.1.1, training an RF classifier using the same Euclid-only features listed in Footnote 7.
+However, this time we labelled the sample based on the classification derived from the versus plot, setting Galactic objects (396) to ‘1‘ and extragalactic objects (1000) to ‘0‘.
+Using the trained model, we assigned a Galactic or extragalactic label to each Euclid CTP, providing a probabilistic classification, , for robust downstream analyses.
 We present the respective confusion matrix in Fig.
 10 and note a recall of 92%, with a limited fall-out of 4%.
-For more details on the feature correlations see Appendix A.
+For more details on the feature correlations, see Appendix A.
 To further validate the RF performance, we examine the galactic probability distribution () for the unique CTP candidates sample.
@@
 6 Redshifts
-The catalogues of CTPs we make available in this paper include spectroscopic redshifts (spec-s), where available, by matching their coordinates to a compilation of publicly available redshifts (Kluge et al.
-2024).
-However, this compilation is rich in duplications, hence the details of the cleaning are described in Sect.
+The catalogues of CTPs we make available in this paper include spectroscopic redshifts (spec-s), where available.
+This was done by matching their coordinates to a compilation of publicly available redshifts (Kluge et al.
+2024, Igo et al.
+in prep.).
+However, this compilation is rich in duplications; hence, the details of the cleaning are described in Sect.
 3.1 of Saxena et al.
@@
 While spec-s provide highly accurate distance measurements, they are inherently limited to a subset of typically brighter objects.
-As a result only 16% of the CTP sample benefits from publicly available spec- coverage (refer to Footnote 8).
+As a result, only 16% of the CTP sample benefits from publicly available spec- coverage (refer to Footnote 9).
 6.1 Photometric redshifts
-To achieve a more complete characterization of the CTPs, photometric redshifts (photo-s) are essential.
+To achieve a more complete characterisation of the CTPs, photometric redshifts (photo-) are essential.
 However, achieving reliable photo-s for AGN remains significantly more challenging than for inactive galaxies, despite notable advances in recent years (see Salvato et al.
 2019, for a review).
-The primary difficulty lies in the overwhelming emission produced by the accretion onto the SMBH, which often masks the spectral features of the host galaxy, crucial for breaking colour-redshift degeneracies.
-The current version of photo-s released for Q1 assumes the sources to be inactive galaxies and does not account for the possible contribution from an active nucleus.
-The absence of AGN SEDs in the template library makes it difficult to accurately constrain their contributions using photometry alone, resulting in miscalculations of the physical parameters for each source derived indirectly.
+The current Q1 photo-s (Euclid Collaboration: Tucci et al.
+2025), derived using Phosphoros (Paltani et al.
+in prep.), are estimated from total fluxes and fine-tuned for inactive galaxies.
+This approach neglects the photometric signatures of active nuclei, whose emission can dominate the observed flux and hide key host galaxy features, crucial for breaking colour-redshift degeneracies (Salvato et al.
+2019).
 Instead, studies have demonstrated that incorporating aperture or pixel-based analyses can significantly enhance the accuracy of such photo- estimates.
 Consequently, to compute reliable photo-s for AGN-dominated objects in a dedicated effort, we utilise PICZL (Roster et al.
-2024), an ML algorithm that significantly improves photo- estimation for AGN by directly predicting redshifts from imaging data, eliminating the need for manual feature extraction or additional surveys.
+2024), an ML algorithm that significantly improves photo- estimation for AGN by directly predicting redshifts from imaging, eliminating the need for manual feature extraction or combination of various surveys.
 Building on its predecessor, Circlez (Saxena et al.
 2024), which analysed aperture-based photometric variations, PICZL advances to pixel-level resolution, leveraging the raw images for improved precision.
-As a result, for sources bright enough to successfully be matched to the LS10 catalogue, we are able to produce reliable photo-s (refer to Footnote 8).
-This is possible since the training sample constructed for PICZL is formed mostly by X-ray selected AGN detected in LS10, enabling us to use the same model.
-As in the original work, the uncertainty of the photo-s in terms of 1 error, increases with the faintness of the sources (see Roster et al.
+As a result, for sources bright enough to successfully match the LS10 catalogue, we are able to produce reliable photo-s (refer to Footnote 9).
+As in the original work, the uncertainty of the photo-s in terms of 1 error increases with the faintness of the sources (see Fig.
+12 and Roster et al.
 2024, for more details).
-For all EDF CTPs of an X-ray source detected in LS10, PICZL provides full redshift posterior distributions (PDZ) derived from de-reddened calibrated images in the , , , and bands, supplemented by , , , and catalogue-based aperture photometry without imposing any S/N cuts.
+This study does not incorporate Euclid photometry in the photo- estimation, as doing so would require a dedicated pipeline, including re-training on Euclid imaging and assembling a spectroscopic AGN training set from the Q1 field.
+Given the small and incomplete nature of such a sample, especially at the faint end, this effort is deferred until DR1 becomes available.
+An additional challenge lies in the heterogeneous photometric coverage across Q1: while Euclid provides uniform measurements in , , , and , these are complemented by ground-based data of varying depth and filter sets, leading to inconsistencies in photo- quality.
+Consequently, to ensure robustness and comparability, we opted to rely on established methods, with PICZL offering high completeness and reliable redshifts for 82% of the X-ray sources detected in LS10101010Future PICZL versions will make use Euclid imaging, possibly combined with LSST (Ivezic et al.
+2019).
+(see Fig.
+13).
+Nonetheless, we recognise the potential of Euclid photometry to enhance our performance, particularly for faint sources and at high redshifts (Graham et al.
+2020).
+As illustrated in the left panel of Fig.
+12, numerous faint sources are dispersed horizontally around , as no strong spectral breaks can be captured by the Legacy filters around this redshift (Roster et al.
+2024).
+Access to Euclid ’s NIR bands , , and will enable more accurate redshift recovery in this range.
+For all EDF CTPs of an X-ray source detected in LS10, PICZL provides full redshift posterior distributions (PDZ), derived from de-reddened calibrated images in the , , , and bands, supplemented by , , , and catalogue-based aperture photometry without imposing any S/N cuts.
 The method is agnostic to the morphological classification of the source, applying equally to extended and point-like objects.
-As output, PICZL provides the dominant mode of the PDZ as a point prediction, along with uncertainties quantified as 1 and 3 confidence intervals.
+As output, PICZL provides the dominant mode of the PDZ as a point prediction, along with uncertainties quantified as 1 confidence intervals.
 Notably, we compute photo-s also for those CTPs matched to LS10, which we believe to be Galactic.
@@
 6.2 Photo-z quality
-To evaluate the reliability of our photo- estimates, we employ standard metrics commonly implemented in the literature, including (a) the accuracy as a measure of the scatter between the prediction and truth values of the sample, defined as
-and (b) the fraction of outliers, , which quantifies the proportion of sources whose redshift estimates deviates significantly from their corresponding spectroscopic redshift.
+To evaluate the reliability of our photo- estimates, we employed standard metrics commonly implemented in the literature, including (a) the accuracy as a measure of the scatter between the prediction and truth values of the sample, which is defined as
+and (b) the fraction of outliers, , that quantifies the proportion of sources whose redshift estimates deviate significantly from their corresponding spectroscopic redshift.
 Specifically, this is defined as the fraction of objects for which
-We evaluated the performance of the photo-s by comparing them to the spec-s, as shown in Fig.
+We evaluated the performance of the photo-s by comparing them to the spec-s, as shown in the middle panel of Fig.
 12.
-For the subsample of sources with ¡ 21.5, which corresponds to the brighter and more reliable sources in our data set, we find that the fraction of outliers is approximately 5%, with a of roughly 0.04 as opposed to and observed for the entire sample.
-Additionally, the catastrophic outliers at are faint, galaxy-dominated objects ( ¿ 24) with large photometric errors in LS10.
-In scientific studies involving photometric redshifts, it is crucial to account for such uncertainties and in turn, photometric redshift algorithms should therefore make sure to provide reliable errors.
-PICZL achieves this effectively as discussed in detail in Roster et al.
-(2024).
-In the right panel of Fig.
-12 this is evidenced by the smaller errors closely tracing the identity line, while larger errors appropriately account for sources of uncertainty.
+For the subsample of bright sources with ¡ 21.5, corresponding to those that have smaller photometric errors within our dataset, we find a fraction of outliers of 5.4%, with a of roughly 0.04 as opposed to and observed for the entire sample.
+The catastrophic outliers at appear as faint ( ¿ 24), galaxy-dominated objects with larger photometric errors in LS10 and highly degenerate PDZs.
+In scientific studies involving photo-s, it is crucial to account for such uncertainties and in turn, photo- algorithms should provide reliable error estimates.
+PICZL achieves this effectively as shown in the right panel of Fig.
+12, with small errors closely following the identity line and larger ones capturing photometric uncertainties.
+Despite occasional higher uncertainties, point estimates remain robust, ensuring that for applications such as luminosity functions, the entire PDZs remain valuable.
 In Fig.
-13, we present the cumulative distribution of CTP band magnitudes along with the respective fractions of sources with spectroscopic and/or and photometric redshifts.
-The figure reveals that the sample is almost complete in redshift up to around .
-For fainter sources, photo- estimation is either not possible due to the limited number of photometric data points and significant errors, or the source, though clearly detected in Euclid, is not visible in LS10 due to its shallower depth.
-In the future, when we have upgraded PICZL to be applied to Euclid images and have integrated data from the Legacy Survey of Space and Time (LSST; Ivezic et al.
-2019), photometric redshifts for AGN in Euclid will dramatically improve.
+13, we present the cumulative distribution of CTP band magnitudes along with the corresponding fractions of sources for which spec- and/or and photo-s from PICZL are available.
+The figure demonstrates that redshift completeness is very high up to around , beyond which the availability of reliable photo- estimates begins to decline.
+In addition to spectroscopic selection effects, this is due to fainter sources which, while clearly detected in Euclid, often lack sufficient optical coverage or exhibit photometric uncertainties too large to yield robust photo- values.
+Future releases of PICZL, adapted to directly process Euclid imaging and incorporating the deeper and more complete photometry expected from the Legacy Survey of Space and Time (LSST; Ivezic et al.
+2019), will markedly enhance photo- performance for fainter AGN.
+We compared our photo- estimates with those from Duncan (2022), who combined ML and SED fitting to derive redshifts for all extragalactic sources in the 8th data release of the Legacy Survey (LS8).
+In EDFN, where LS10 photometry corresponds roughly to that of LS8, PICZL achieves a slight improvement in (from 0.10 to 0.08) and a notable reduction in (from 25% to 14%), without applying magnitude cuts.
+In the two additional fields covered by LS10-South (Zenteno et al.
+2025), where PICZL also benefits from deeper WISE and added i-band photometry, the improvements are even more pronounced, with improving from 0.09 to 0.05, and dropping from 23% to 12%.
+Likewise, while the photo-s from Zhou et al.
+(2021, 2023), computed with LS10 photometry, perform extremely well for inactive galaxies, a trend similar to Duncan (2022) is observed for AGN, where the accuracy is known to decline111111with reaching up to 33% in some subsets (priv.
+comm.).
+For high-redshift ( ¿ 3) X-ray selected AGN, we compared our estimates to those of Pouliasis et al.
+(2025), computed using LS10 supplemented with near-infrared photometry from the VISTA Hemisphere Survey (VHS, McMahon et al.
+2013).
+Of the 45 sources in common, only two have spec- for which both photo- agree, preventing any meaningful additional comparison.
+Instead, we can assess our photo- quality for X-ray selected AGN at the faint end by benchmarking them against studies that employ significantly deeper and broader photometric datasets.
 7 Properties of the CTP sample
@@
 This refined selection yields a total of 9294 sources from the initial 11 286 entries having match_flag=1.
-Figure 14 illustrates the magnitude distribution of our CTPs, grouped by the match between EDFs and the corresponding X-ray survey (4XMM, CSC2, and eROSITA DR1, respectively).
-The 4XMM and CSC 2.0 distributions show their ability to detect fainter X-ray sources, which correlates with the identification of fainter optical CTPs.
+Building on this approach, we note that a similar comparison has already been presented in Roster et al.
+(2024), where PICZL photo-s were evaluated against the XMM-SERVS catalogues from Chen et al.
+(2018) and Ni et al.
+(2021).
+Here, we extend this strategy by comparing our photo- results with the 7 Ms Source Catalog from the Deep Field-South (CDFS, Luo et al.
+2017).
+This catalogue is particularly well-suited for faint-end validation, as it includes photo-s derived using high-quality multi-band data, largely informed by Hsu et al.
+(2014).
+Utilising up to 38 intermediate and broad bands spanning the ultraviolet (GALEX, Martin et al.
+2005) to the mid-infrared (Spitzer/IRAC, Cardamone et al.
+2010), their work has resulted in one of the most comprehensive SED-fitting datasets for faint X-ray-selected sources to date.
+To ensure a proper one-to-one comparison between sources, we match our CTP catalogue to that of Luo et al.
+(2017) using the respective CTP coordinates.
+We recover the same CTPs across the full sample, with 97.5% of sources showing offsets below 08, with the largest observed separation being 18.
+Indicating robust matches, this consistency highlights the overall quality of our CTPs that can be attributed to the effectiveness of the Euclid-based prior used with NWAY, providing a solid foundation for future applications.
+For the 400 sources with spec-s matched between this work and Luo et al.
+(2017), we compare the two datasets by evaluating and as a function of -band magnitude from the Wide Field Imager (WFI; Giavalisco et al.
+2004).
+As depicted in Fig.
+23, PICZL achieves a lower fraction of outliers than Luo et al.
+(2017) for 22.3, though at larger due to using only four shallower optical and WISE mid-IR bands.
+However, even up to PICZL extrapolation results in outlier fractions of , which is lower compared to, for example, eROSITA / eFEDS (Salvato et al.
+2022), where the fraction is %, despite using many more and deeper bands.
+Given the overall quality of Luo et al.
+(2017), we decided to extend redshift completeness in this work by incorporating 378 unique photo-s from the 7 Ms Source Catalog.
+Figure 14 illustrates the magnitude distribution of our CTPs, grouped by the match between EDFs and the corresponding X-ray survey (4XMM DR14, CSC 2.0, and eROSITA DR1, respectively).
+The 4XMM and CSC2 distributions show their ability to detect fainter X-ray sources, which correlates with the identification of fainter optical CTPs.
 In contrast, the eROSITA DR1 distributions exhibit brighter median magnitudes, consistent with the shallowness of the survey (see Fig.
@@
 Additionally, the figure suggests a possible bimodal distribution at the faint end in the optical band for both 4XMM and CSC 2.0, at least in EDF-N and EDF-S.
-The majority of ources in the second, faint peak of the distribution show X-ray fluxes , with no obvious drop off in p_any, and the distribution will be studied further in other papers.
+The majority of sources in the second, faint peak of the distribution show X-ray fluxes , with no obvious drop off in p_any, and the distribution will be studied further in other papers.
 7.1 Colour-redshift relation
-Redshifts, spectroscopic and photometric, are available for 79% of the CTPs, allowing us to analyse their Euclid colours as function of redshift, as shown in Fig.
+Redshifts, spectroscopic and photometric, are available for 81% of the CTPs, allowing us to analyse their Euclid colours as a function of redshift, as shown in Fig.
 15.
@@
 2021).
-As iterated before, the Euclid filters alone lack the fidelity to catch variations in the SED of sources, a caveat that we will be able to overcome in combination with upcoming surveys, especially LSST.
+As iterated before, Euclid filters alone lack the fidelity to catch variations in the SED of sources, a caveat that we will be able to overcome in combination with upcoming surveys, especially LSST.
 7.2 X-ray luminosities
-For the rest of the analysis, we restrict the CTP sample to those 7345 extragalactic sources with a soft X-ray flux S/N 2 in order to reduce potential spurious X-ray detection.
+For the rest of the analysis, we restrict the CTP sample to those 7345 extragalactic sources with a soft X-ray flux S/Nx 2 in order to reduce potential spurious X-ray detections.
 In Fig.
 16, we observe a clear distinction between pointed observations taken with XMM-Newton and Chandra at various depths versus the homogenous all-sky coverage provided by eROSITA.
-The complete coverage of eROSITA results in a tighter correlation between luminosity and redshift, despite the presence of a few points away from the general trend, which have low S/N (thus indicating potential spurious sources).
-Another reason for their offset from the general trend could be an incorrect photometric redshift.
-On the other hand, XMM-Newton and Chandra catalogues include a mixture of pointed observations resulting in a less smooth distribution between redshift and luminosity.
-In all three panels, we indicate with open circle sources with low S/Nr, potentially corresponding to those highlighted red in the left panel of Fig.
-12, representing examples of increased redshift uncertainties.
-Notably, most of the results shown, focus on objects with redshifts of .
+The complete coverage of eROSITA results in a tighter correlation between luminosity and redshift, despite the presence of a few points away from the general trend, with S/N.
+Another reason for their offset from the general trend could be erroneous photo-.
+On the other hand, XMM-Newton and Chandra catalogues include a mixture of pointed observations, resulting in a less smooth distribution between redshift and luminosity.
+In all three panels, we indicate with open circles sources with low S/Nr, presumably representing examples of increased redshift uncertainties.
+Notably, most of the results shown focus on objects with redshifts of .
 Sources with higher redshifts, while present, should be approached with caution due to potential inaccuracies.
@@
 2013).
-Applying the criterion of , we classify 6243 out of the 6295 sources with computed X-ray luminosities as AGN.
-A more inclusive approach is presented in Georgakakis et al.
-(2007) for example, where X-ray detected starburst galaxy candidates, though more likely in case of Chandra as opposed to eROSITA (Kyritsis et al.
+According to Georgakakis et al.
+(2007), for example, X-ray detected starburst galaxy candidates, though more likely in case of Chandra as opposed to eROSITA (Kyritsis et al.
 2025), can represent about 20% of the X-ray source population in the luminosity interval .
-This fraction is also likely to be an upper limit because some of the galaxy candidates may turn out to be low–luminosity AGN, such that distinguishing between these sources can be challenging.
-Using our sources from XMM-Newton, where both soft and hard (2 – 12 keV)999We assume that at zero order, the luminosities between (2 – 10) and (2 – 12) keV are approximately the same.
-band fluxes are available, we show that, to retain all sources in this parent catalogue that are classified as AGN using the hard band, a soft-band luminosity threshold of is required (refer to Fig.
+This fraction is also likely to be an upper limit because some of the galaxy candidates may turn out to be low-luminosity AGN, such that distinguishing between these sources can be challenging.
+Consequently, with sources from XMM-Newton, where both soft and hard (2–12 keV)121212We assume that at zero order, the luminosities between (2–10) and (2–12) keV are approximately the same.
+band fluxes are available, we show that, to retain all sources in this parent catalogue that are classified as AGN using the hard band limit, a soft-band luminosity threshold of is required (refer to Fig.
 17).
@@
 Applying this more inclusive threshold defined in Fig.
-17 to our extragalactic CTP sample, 6292 of the 6295 sources that we were able to compute X-ray luminosities for, meet the AGN definition criterion.
+17 to our extragalactic CTP sample, all of the 6653 (91%) sources that we were able to compute X-ray luminosities for, meet the AGN definition criterion, 49 of which have .
 7.3 AGN type
@@
 We add the —X/O— = 1 lines (Maccacaro et al.
-1988), which highlight the typical AGN locus (see also e.g., Salvato et al.
+1988), which confine the typical AGN locus (see also e.g., Salvato et al.
 2011; Civano et al.
@@
 For we find .
-To further assess the type of our sources, we colour-code the plots in the bottom row of Fig.
-18 by the X-ray luminosity and the top row by hardness ratio (Park et al.
-2006; Cappelluti et al.
+To further assess the AGN type of our CTPs, we colour-code the plots in the bottom row of Fig.
+18 by the X-ray luminosity and the top row by hardness ratio (Hasinger et al.
+2005; Cappelluti et al.
 2007; Salvato et al.
-2008), defined as
-where a negative HR value indicates (type I) unobscured sources, whereas a positive HR value suggests harder, likely obscured (type II) sources.
-Therefore, we include fluxes from the hard X-ray band where available.
-Specifically, we consider fluxes from the 2.0 – 4.5 keV range for 4XMM DR14, the 2.0 – 7.0 keV range for Chandra, and the 2.3 – 5.0 keV energy range for eROSITA DR1.
-This ensures as much consistency as possible across the different surveys, allowing us to utilise similar energy bands for comparison and analysis.
-Since the markers in Fig.
-18 are plotted in order of increasing hardness ratio, many sources in the background are being covered up and predominantly fall within the X/O boundaries, classifying them as bright, unobscured type I sources, with different populations visible for every X-ray survey.
-This trend is most pronounced for eROSITA, where the distribution is more concentrated within the boundaries.
-A similar pattern is observed for XMM-Newton, though it includes a population of very soft sources lying below the X/O boundaries, potentially corresponding to starburst galaxies.
-For Chandra, the general trend also holds; however, there is an additional presence of very hard sources overlapping with very soft ones below the X/O boundary, particularly in the bright optical and faint X-ray regime.
-The majority of these outliers are sources with S/N 3 in the soft X-ray band, suggesting that at least some of them may be spurious detections.
+2009), defined as
+where soft corresponds to the 0.5–2.0 keV detection band, and hard refers to 2.0–12.0 keV PN thin for 4XMM DR14, 2.0–7.0 keV ACIS-I for Chandra, and 2.0–5.0 keV for eROSITA DR1 respectively, to account for each instrument’s flux sensitivity.
+This ensures as much consistency as possible across the different surveys, to make the different instruments somewhat comparable.
+A negative HR value indicates soft, unobscured sources (type I), and a positive HR value suggests harder, likely obscured sources (type II).
+We examine the HR distributions for all three telescopes to assess differences in their observed AGN populations.
+However, a non-negligible fraction of sources lack hard-band detections: approximately 47.6% for eROSITA, 14.5% for XMM-Newton, and 7.3% for .
+These non-detections, especially in the case of eROSITA, naturally shift the medians of the true HR distributions and introduce selection biases.
+As shown in Fig.
+24, we find that the eROSITA HR distribution median is , while that of and (XMM-Newton) show a similar tendency towards harder values with medians of .
+These HR medians can then be used to compute ‘effective’ photon indices , which we find to be for eROSITA, for 131313Adopting CHANDRA-Cycle 11 as a representative benchmark, as it reflects an intermediate instrumental state within the 20-year span of observations included in CSC 2.0, accounting for the effects of detector degradation over time.
+and for XMM-Newton, assuming negligible column density.
+As expected, the greater sensitivity of XMM-Newton and at higher energies allows them to detect harder and possibly mildly obscured AGNs that are missed by eROSITA.
+Conversely, eROSITA tends to detect brighter AGN which exhibit softer X-ray spectra.
+As the markers in Fig.
+18 are plotted in order of increasing HR, many softer sources in the background are being covered up and predominantly fall within the X/O boundaries, classifying them as AGN, with differently weighted populations visible for every X-ray survey.
+As observed for XMM-Newton and , this includes a population of sources with lying below the X/O boundaries, potentially corresponding to starburst galaxies, low-luminosity AGN, or obscured AGN.
+However, the majority of these sources show S/Nx 3 in the soft X-ray band, suggesting that a few of these may be spurious detections.
 It is also plausible that these sources are intrinsically variable, suggesting that they were observed in different states of appearance.
@@
 8 Release of the catalogue
-The catalogue containing the CTPs to the point-like X-ray sources in the EDFs and their properties are available through CDS/Vizier101010https://vizier.cds.unistra.fr.
-A detailed description of the columns and their content is provided in Appendix B.
-The catalogue includes unique identifiers for all relevant surveys as well as the basic X-ray and Euclid properties (columns 1 – 13).
+The catalogue containing the CTP properties of the EDF point-like X-ray sources is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
+A detailed description of the columns and their content is provided in Appendix D.
+The catalogue includes unique identifiers for all relevant surveys as well as the basic X-ray and Euclid properties (columns 1–14).
 For the full list of available columns, please refer to the original catalogues introduced in Sect.
-3 and e.g., Euclid Collaboration: McCracken et al.
+3 and, for example, Euclid Collaboration: McCracken et al.
 (2025); Euclid Collaboration: Polenta et al.
@@
 (2025).
-Columns 14 – 24 present the results of the CTP association, followed by key parameters derived from NWAY.
-Photometric data from LS10 are reported in columns 25 – 28.
-Spectroscopic information, when available, is included in columns 29 – 33.
-Lastly, the photometric redshift parameters from the PICZL algorithm, as well as quantities computed with such, are listed in columns 34 – 40.
+Columns 15–24 present the results of the CTP association, followed by key parameters derived from NWAY.
+Photometric data from LS10 are reported in columns 25–27.
+Spectroscopic information, when available, is included in columns 28–32.
+Lastly, the photometric redshift parameters from the PICZL algorithm, as well as quantities computed with such, are listed in columns 33–39.
 9 Conclusions and outlook
-In this paper we have presented the procedure adopted for the identification, classification, and study of the Euclid CTPs to the point-like sources selected from XMM-Newton, Chandra, and eROSITA X-ray, detected in Q1.
+In this paper, we have presented the procedure adopted for the identification, classification, and study of the Euclid CTPs of the point-like sources selected from XMM-Newton, Chandra, and eROSITA X-ray, detected in Q1.
 We summarise the most important results in the following.
 Utilising NWAY (Salvato et al.
-2018), supplemented by a prior on X-ray emission, based on properties of a 4XMM DR11, and CSC 2.0 training sample, we achieved highly accurate CTP identifications relying on Euclid photometry only (see Sect.
+2018) supplemented by a prior on X-ray emission based on properties of a 4XMM DR11 and CSC 2.0 training sample, we achieved highly accurate CTP identifications while relying on Euclid photometry only (see Sect.
 4).
-Randomising the X-ray position and repeating the same procedure for the association, we quantify for each X-ray source the probability of chance associations.
-Users will be able to create subsamples of CTPs of desired purity and completeness by applying more restrictive or inclusive p_any (the probability of an X-ray source having a CTP) cuts (see Fig.
+By randomising the X-ray position and repeating the same procedure for the association, we quantified for each X-ray source the probability of chance associations.
+Users will be able to create subsamples of CTPs of desired purity and completeness by applying more restrictive or inclusive p_any cuts ( probability of an X-ray source having a CTP; see Fig.
 7).
-Altogether, we identify 12 645 CTPs, with only about 10% of the sources having more than a single possible CTP, mostly in the 4XMM and eROSITA surveys, due to their larger positional uncertainties, compared to Chandra.
-We attribute this performance to the availability of extensive X-ray detected samples with secure CTPs, and deep, homogenised multiwavelength photometry spanning optical to MIR wavelengths, enabling robust SED construction for training.
-Further enhancements in expanding to broader multiwavelength coverage will improve the quality of CTP identification, particularly with Euclid DR1 in mind.
-We trained a RF algorithm to classify objects as Galactic or extragalactic using solely Euclid photometry in Sect.
+Altogether, we have identified 12 645 CTPs, with only about 10% of the sources having more than a single possible CTP.
+These are mostly in the XMM-Newton and eROSITA surveys, due to their larger positional uncertainties compared to Chandra.
+We attribute this performance to the availability of extensive X-ray detected samples with secure CTPs and deep, homogenised multiwavelength photometry spanning optical to MIR wavelengths enabling robust SED construction for training.
+Further expansion to broader multiwavelength coverage will improve the quality of CTP identification, particularly for Euclid DR1.
+We trained an RF algorithm to classify objects as Galactic or extragalactic using solely Euclid photometry in Sect.
 5.
-We previously trained on a sample of secure Galactic/extragalactic sources detected in LS10, using a over W1 thresholds, as presented in Salvato et al.
+We previously trained another RF on a sample of secure Galactic and extragalactic sources detected in LS10, using over W1 thresholds, as presented in Salvato et al.
 (2022), effectively breaking the degeneracy in classification that arises when relying on Euclid photometry alone (see Euclid Collaboration: Bisigello et al.
 2024).
-We validate these classifications by comparing to Euclid Collaboration: Matamoro Zatarain et al.
+We validated these classifications by comparing to Euclid Collaboration: Matamoro Zatarain et al.
 (2025).
 We also added spectroscopic redshifts for 16% of the CTP sample using publicly available data.
-For the 76% of the sample detected in LS10, we computed photometric redshifts using PICZL (Roster et al.
-2024), reaching up to , although at these high redshifts the values are associated with large uncertainties.
-With all these redshifts, we investigated the colour-redshift relation of our sample in addition to computing their X-ray luminosities (refer to Sect.
+For the 82% of the sample detected in LS10, we computed photo- using PICZL (Roster et al.
+2024), reaching up to .
+However at these high redshifts, the values are associated with large uncertainties.
+With all of these redshifts, we investigated the colour-redshift relation of our sample in addition to computing their X-ray luminosities (refer to Sect.
 6).
-We find that almost all of our CTPs pass the threshold to be considered an AGN, with most of them appearing as soft, bright, and unobscured type I sources.
+We find that all of our CTPs pass our soft X-ray luminosity threshold to be considered an AGN, with more of them appearing as bright type I sources rather than type II.
 As the primary outcome of this paper, we present the CTP catalogue, which includes a wealth of information such as IDs, basic X-ray properties, NWAY outputs, classification probabilities, and redshift data (see Sect.
@@
 By providing probabilities for all applicable sample characteristics, we ensure users have the flexibility to make informed selections tailored to their specific scientific goals, without being constrained by predefined thresholds or assumptions.
-The RF models used for the identification of the CTPs and their classification into Galatic/extragalactic, can be applied blindly to the entire Q1 catalogue.
-For each source, they provide the probability of being an X-ray emitter, as well as an extragalactic source, despite the fact of not having been detected by any of the three X-ray surveys yet.
-Consequently, as illustrated in Figure Fig.
-19, limiting the selection of Q1 sources to those with and , while excluding Euclid sources classified as CTPs in this study, yields a final sample of 135 303 objects.
-Notably, approximately 400 of these have been successfully cross-matched geometrically with the little red dots sample introduced by Euclid Collaboration: Bisigello et al.
-(2025), predominantly occupying the bright end of the magnitude distribution.
-A forthcoming paper will explore the multi-wavelength nature of likely extragalactic high sources that remain undetected in current X-ray surveys.
+The RF models used for the identification of the CTPs and their classification as Galactic or extragalactic, can be applied blindly to the entire Q1 catalogue.
+For each source, they provide the probability of being an X-ray emitter as well as an extragalactic source, despite not having been detected by any of the three X-ray surveys yet.
+Consequently, as illustrated in Fig.
+19, limiting the selection of Q1 sources to those with and while excluding Euclid sources classified as CTPs in this study yields a final sample of 135 303 objects.
+Notably, 73 of these objects have been successfully cross-matched geometrically with the little red dots (LRDs) sample introduced by Euclid Collaboration: Bisigello et al.
+(2025), where the matches predominantly occupy the bright end of the LRD magnitude distribution.
+A forthcoming paper will explore the multi-wavelength nature of the likely extragalactic high sources that remain undetected in current X-ray surveys.
 Acknowledgements.
@@
 - Cappelluti et al. (2007) Cappelluti, N., Hasinger, G., Brusa, M., et al. 2007, ApJS, 172, 341
+- Cardamone et al. (2010) Cardamone, C. N., van Dokkum, P. G., Urry, C. M., et al. 2010, ApJS, 189, 270
 - Cavuoti et al. (2014) Cavuoti, S., Brescia, M., D’Abrusco, R., Longo, G., & Paolillo, M. 2014, MNRAS, 437, 968
+- Chen et al. (2018) Chen, C.-T. J., Brandt, W. N., Luo, B., et al. 2018, MNRAS, 478, 2132–2163
 - Chen et al. (2022) Chen, H., Lundberg, S. M., & Lee, S.-I. 2022, Nature Communications, 13
@@
 - Cooper et al. (2023) Cooper, N., Dainotti, M. G., Narendra, A., Liodakis, I., & Bogdan, M. 2023, MNRAS, 525, 1731
-- Cuillandre et al. (2024) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2024, A&A, accepted, arXiv:2405.13501
+- Cuillandre et al. (2025) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2025, A&A, 697, A11
 - Daoutis et al. (2023) Daoutis, C., Kyritsis, E., Kouroumpatzakis, K., & Zezas, A. 2023, A&A, 679, A76
@@
 - Donley et al. (2012) Donley, J. L., Koekemoer, A. M., Brusa, M., et al. 2012, ApJ, 748, 142
+- Duncan (2022) Duncan, K. J. 2022, MNRAS, 512, 3662
 - Ebrero et al. (2009) Ebrero, J., Carrera, F. J., Page, M. J., et al. 2009, A&A, 493, 55
 - Elvis et al. (1994) Elvis, M., Wilkes, B. J., McDowell, J. C., et al. 1994, ApJS, 95, 1
-- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted
+- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15302
 - Euclid Collaboration: Bisigello et al. (2024) Euclid Collaboration: Bisigello, L., Massimo, M., Tortora, C., et al. 2024, A&A, 691, A1
-- Euclid Collaboration: Bisigello et al. (2025) Euclid Collaboration: Bisigello, L., Rodighiero, G., Fotopoulou, S., et al. 2025, A&A, submitted
-- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
+- Euclid Collaboration: Bisigello et al.
+(2025) Euclid Collaboration: Bisigello, L., Rodighiero, G., Fotopoulou, S., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554537, arXiv:2503.15323
+- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
 - Euclid Collaboration: Desprez et al. (2020) Euclid Collaboration: Desprez, G., Paltani, S., Coupon, J., et al. 2020, A&A, 644, A31
-- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
+- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
 - Euclid Collaboration: Lusso et al. (2024) Euclid Collaboration: Lusso, E., Fotopoulou, S., Selwood, M., et al. 2024, A&A, 685, A108
-- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
-- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H., Benson, K., et al. 2025, A&A, submitted
-- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
-- Euclid Collaboration: Polenta et al. (2025) Euclid Collaboration: Polenta, G., Frailis, M., Alavi, A., et al. 2025, A&A, submitted
-- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
+- Euclid Collaboration: Matamoro Zatarain et al.
+(2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554619, arXiv:2503.15320
+- Euclid Collaboration: McCracken et al.
+(2025) Euclid Collaboration: McCracken, H.
+J., Benson, K., Dolding, C., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554594, arXiv:2503.15303
+- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
+- Euclid Collaboration: Polenta et al.
+(2025) Euclid Collaboration: Polenta, G., Frailis, M., Alavi, A., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554657, arXiv:2503.15304
+- Euclid Collaboration: Romelli et al.
+(2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554586, arXiv:2503.15305
 - Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
 - Euclid Collaboration: Selwood et al. (2025) Euclid Collaboration: Selwood, M., Fotopoulou, S., Bremer, M. N., et al. 2025, A&A, 693, A250
+- Euclid Collaboration: Tucci et al.
+(2025) Euclid Collaboration: Tucci, M., Paltani, S., Hartley, W.
+G., et al.
+2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554588, arXiv:2503.15306
 - Evans et al. (2024) Evans, I. N., Evans, J. D., Martínez-Galarza, J. R., et al. 2024, ApJS, 274, 22
@@
 - Georgakakis et al. (2007) Georgakakis, A., Rowan-Robinson, M., Babbedge, T. S. R., & Georgantopoulos, I. 2007, MNRAS, 377, 203
+- Giavalisco et al. (2004) Giavalisco, M., Ferguson, H. C., Koekemoer, A. M., et al. 2004, ApJ, 600, L93
+- Graham et al. (2020) Graham, M. L., Connolly, A. J., Wang, W., et al. 2020, AJ, 159, 258
 - Green et al. (2024) Green, K., Elmer, E., Maltby, D. T., et al. 2024, MNRAS, 531, 2551
@@
 (2022) Maciaszek, T., Ealet, A., Gillard, W., et al.
-2022, in SPIE, Vol.
+2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol.
 12180, Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, ed.
@@
 - Marchesi et al. (2016) Marchesi, S., Lanzuisi, G., Civano, F., et al. 2016, ApJ, 830, 100
+- Martin et al. (2005) Martin, D. C., Fanson, J., Schiminovich, D., et al. 2005, ApJ, 619, L1
+- McMahon et al. (2013) McMahon, R. G., Banerji, M., Gonzalez, E., et al. 2013, The Messenger, 154, 35
 - Mechbal et al. (2024) Mechbal, S., Ackermann, M., & Kowalski, M. 2024, A&A, 685, A107
@@
 - Naylor et al. (2013) Naylor, T., Broos, P. S., & Feigelson, E. D. 2013, ApJS, 209, 30
+- Ni et al. (2021) Ni, Q., Brandt, W. N., Chen, C.-T., et al. 2021, ApJS, 256, 21
 - Oh et al. (2018) Oh, K., Koss, M., Markwardt, C. B., et al. 2018, ApJS, 235, 4
@@
 - Palanque-Delabrouille et al. (2016) Palanque-Delabrouille, N., Magneville, C., Yèche, C., et al. 2016, A&A, 587, A41
-- Park et al. (2006) Park, T., Kashyap, V. L., Siemiginowska, A., et al. 2006, ApJ, 652, 610
 - Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
@@
 - Pineau et al. (2017) Pineau, F. X., Derriere, S., Motch, C., et al. 2017, A&A, 597, A89
+- Pouliasis et al. (2025) Pouliasis, E., Ruiz, A., Georgantopoulos, I., et al. 2025, A&A, 697, A123
 - Pović et al. (2012) Pović , M., Sánchez-Portal, M., García, A. M. P., et al. 2012, A&A, 541, A118
@@
 - Salvato et al. (2018) Salvato, M., Buchner, J., Budavári, T., et al. 2018, MNRAS, 473, 4937
-- Salvato et al. (2008) Salvato, M., Hasinger, G., Ilbert, O., et al. 2008, ApJ, 690, 1250
+- Salvato et al. (2009) Salvato, M., Hasinger, G., Ilbert, O., et al. 2009, ApJ, 690, 1250
 - Salvato et al. (2011) Salvato, M., Ilbert, O., Hasinger, G., et al. 2011, ApJ, 742, 61
@@
 - Temple et al. (2021) Temple, M. J., Hewett, P. C., & Banerji, M. 2021, MNRAS, 508, 737
-- Traulsen et al. (2020) Traulsen, I., Schwope, A. D., Lamer, G., et al. 2020, A&A, 641, A137
 - Urry & Padovani (1995) Urry, C. M. & Padovani, P. 1995, PASP, 107, 803
@@
 - Xue et al. (2011b) Xue, Y. Q., Luo, B., Brandt, W. N., et al. 2011b, ApJS, 195, 10
+- Zenteno et al. (2025) Zenteno, A., Kluge, M., Kharkrang, R., et al. 2025, A&A, 698, A171
 - Zeraatgari et al. (2024) Zeraatgari, F. Z., Hafezianzadeh, F., Zhang, Y., et al. 2024, MNRAS, 527, 4677
+- Zhou et al. (2023) Zhou, R., Ferraro, S., White, M., et al. 2023, J. Cosmology Astropart. Phys., 2023, 097
+- Zhou et al. (2021) Zhou, R., Newman, J. A., Mao, Y.-Y., et al. 2021, MNRAS, 501, 3309
 - Zou et al. (2017) Zou, H., Zhang, T., Zhou, Z., et al. 2017, ApJ, 153
@@
 This not only hinders immediate interpretability but also complicates post-analyses.
-The ML branch investigating the role each feature plays in the model’s predictions is commonly referred to as ’feature importance’.
+The ML branch investigating the role each feature plays in the model’s predictions is commonly referred to as ‘feature importance’.
 To this end, we compute SHapley Additive exPlanations (SHAP; Lundberg & Lee 2017; Rozemberczki et al.
@@
 2022).
-Appendix B Column description of released CTP catalogue
+Appendix B Photometric redshift comparison
+Appendix C HR to effective photon index
+Appendix D Column description of released CTP catalogue
 1.
+Xray_EUCLID_ID: Concatenation of X-ray_ID and EUCLID_ID (unique)
+2.
 Xray_ID: X-ray source identifier (unique)
-2.
+3.
 Xray_RA: J2000 Right Ascension of the X-ray source in the parent catalogue (deg)
-3.
+4.
 Xray_DEC: J2000 Declination of the X-ray source in the parent catalogue (deg)
-4.
+5.
 Xray_CAT: X-ray source catalogue
-5.
+6.
 Xray_pos_err: Positional uncertainty of X-ray source (arcsec)
-6.
+7.
 Fx: Mean 0.5–2 keV X-ray flux from all detections of the X-ray source ()
-7.
+8.
 EFx: Error on the mean 0.5–2 keV X-ray flux from all detections of the X-ray source ()
-8.
+9.
 S_Nx: Signal-to-noise ratio for Fx/EFx
-9.
-Fx_hard: Mean hard (4XMM DR14: 2 – 4.5 keV, CSC 2.0: 2 – 7 keV and eROSITA DR1: 2–5 keV) X-ray flux from all detections of the X-ray source ()
 10.
+HR: Count rate derived hardness ratio considering soft and hard (4XMM DR14: , CSC 2.0: , and eROSITA: ) bands.
+11.
+EDF: Euclid deep field
+12.
 EUCLID_ID: EUCLID source identifier (unique)
-11.
+13.
 EUCLID_RA: J2000 Right Ascension of the counterpart in Q1 catalogue (deg)
-12.
+14.
 EUCLID_DEC: J2000 Declination of the counterpart in Euclid catalogue (deg)
-13.
+15.
 EUCLID_pos_err: Positional uncertainty of Euclid source (arcsec)
-14.
-Xray_EUCLID_ID: Concatenation of X-ray_ID and EUCLID_ID (unique)
-15.
-EUCLID_Xray_proba: Probability to be an X-ray emitter based on the RF
 16.
+EUCLID_P_Xray: Probability to be an X-ray emitter based on the RF
+17.
+EUCLID_P_Gal: Probability of being a Galactic rather than extragalactic source based on the RF
+18.
 Separation_EUCLID_Xray: Distance between Euclid and X-ray source (arcsec)
-17.
+19.
 dist_bayesfactor: Logarithm of ratio between prior and posterior, from separation, positional error, and number density (see Appendix in Salvato et al. 2018)
-18.
-dist_post: Distance posterior probability comparing this association vs. no association (see Appendix in Salvato et al. 2018)
-19.
+20.
+dist_post: Distance posterior probability comparing this association versus no association (see Appendix in Salvato et al. 2018)
+21.
 bias_EUCLID_Xray_proba: Probability weighting introduced by the Euclid-based prior (1 indicates no change)
-20.
+22.
 p_single: Same as dist_post, but weighted by the prior (see Appendix in Salvato et al. 2018)
-21.
+23.
 p_any: Probability that there is a counterpart in LS10 for each X-ray entry (see Appendix in Salvato et al. 2018)
-22.
+24.
 p_i: Relative probability of the X-ray/Euclid match (see Appendix in Salvato et al. 2018)
-23.
+25.
 match_flag: 1 for the most probable match; 2 for almost as good solutions (p_i/p_ibest ¿ 0.5)
-24.
-EDF: Euclid deep field
-25.
+26.
 LS10_FULLID: Unique identifier of the LS10 source matched to the Euclid CTP (concatenation of RELEASE, BRICKID, and OBJID)
-26.
+27.
 LS10_RA: J2000 Right Ascension of the LS10 source (deg)
-27.
+28.
 LS10_DEC: J2000 Declination of the LS10 source (deg)
-28.
-Gal_proba: Probability of being a Galactic rather than extragalactic source based on RF
 29.
@@
 33.
-Cat_spec: Parent catalogue of the spec- match
+Cat_spec: Parent catalogue of the spec-
 34.
@@
 35.
-phz_1sl: Photo- lower 1 sigma error
+phz_1sl: PICZL lower 1 sigma error
 36.
-phz_1su: Photo- upper 1 sigma error
+phz_1su: PICZL upper 1 sigma error
 37.
-phz_3sl: Photo- lower 3 sigma error
+phz_Luo17: Photometric redshift (see Luo et al. 2017)
 38.
-phz_3su: Photo- upper 3 sigma error
+z_final: Spec- when available; otherwise Luo et al. (2017) and PICZL photo-
 39.
-z_final: Spec- when available; otherwise, photo-
-40.
-Lx: X-ray luminosity, assuming FlatLambdaCDM from astropy.cosmology
+Lx: Rest-frame X-ray luminosity, assuming FlatLambdaCDM from astropy.cosmology